### v1

# Ergodicity for locally monotone stochastic evolution equations with Lévy noise

Abstract.
We establish general conditions for stochastic evolution equations with locally monotone drift and degenerate additive Lévy noise in variational formulation resulting in the existence of a unique invariant measure for the associated weakly ergodic Markovian Feller semigroup. We prove improved moment estimates for the solutions and the so-called $e$-property of the semigroup. Examples include the stochastic incompressible 2D Navier-Stokes equations, shear thickening stochastic power-law fluid equations, the stochastic heat equation, as well as, stochastic semilinear equations such as the 1D stochastic Burgers equation.
Key words and phrases:
Additive Gaussian noise; additive Poisson noise; ergodic Markovian Feller semigroup; $e$-property; locally monotone drift stochastic evolution equation with degenerate Lévy noise; stochastic Burgers equation; stochastic incompressible 2D Navier-Stokes equations; stochastic power law fluids; stochastic semilinear equations; unique invariant probability measure
1991 Mathematics Subject Classification:
35K65; 35R60; 37A25; 37L40; 47D07; 60G51; 60H15; 76D05
1. Introduction
In this work, we study ergodicity of the Markovian dynamics associated to a general class of stochastic partial differential equations (SPDEs) on a separable Hilbert space, that is, locally monotone drift stochastic evolution equations with additive Lévy noise, which is the sum of a spatially degenerate Wiener process with possibly infinitely many modes and a degenerate compensated Poisson process with small jumps and finite second moments, which is independent from the Wiener process. Even if the degeneracy of the noise could be seen as a drawback, it is actually this contribution’s strength and main novelty. The main reason is the absence of robust techniques for proving ergodicity for those SPDEs with degenerate noise that are not strongly dissipative and thus exponentially ergodic as in [8].
For non-degenerate noise, more precisely, for noise with minimal non-degeneracy assumptions on the Fourier modes of the spatial structure, the situation is entirely different. We would like to point out that this kind of minimal type of non-degeneracy is sometimes called “degenerate” (as opposed to space-time white noise) or “mildly degenerate” in the literature. In this work, we reserve the terminology “degenerate” for noise which may be zero, finite-dimensional, or spatially regular. Following this terminology, for non-degenerate noise, the main approach derives from the notion of the strong Feller property of the semigroup, combined with irreducibility, which implies uniqueness of the invariant probability measure and therefore ergodicity of the semigroup. This technique works well for both additive and multiplicative noises [20]. It has been refined by Hairer and Mattingly toward the notion of the so-called asymptotic strong Feller property [41, 40], which they used to prove the existence and uniqueness of an invariant measure for the stochastic incompressible 2D Navier-Stokes equations with additive Gaussian forcing [40]. We note that their method yields convergence of the ergodic semigroup in total variation distance, whereas the method used by us yields so-called weak-mean ergodicity, where the time-average of the dual semigroup converges weakly in the sense of probability measures. Our result relies on the ergodicity result of Komorowski, Peszat and Szarek [46], which is a refinement of the so-called lower bound technique of Lasota and Szarek [52], where the main ingredient is the so-called $e$-property of the semigroup, a type of uniform equi-continuity on bounded Lipschitz functions, combined with a asymptotic uniform lower bound for the time-average of the semigroup, which does not require tightness, as the well-known Krylov-Bogoliubov method.
For strongly dissipative monotone drift SPDEs with degenerate Gaussian noise, ergodicity has been obtained, among many other contributions, in [19, 8, 53, 20, 18, 27, 25, 17, 59, 26]. See [76] for a survey article on the topic of ergodicity and Kolmogorov operators for SPDEs. For singular drift monotone SPDEs with degenerate Gaussian noise, ergodicity has been obtained in [34, 28, 35, 73, 57, 78, 29, 10, 54]. For SPDEs with jump noise, ergodicity has for instance been discussed in [31, 66]. See [72] for a recent result on ergodicity of the stochastic Allen-Cahn equation with logarithmic potential.
For the stochastic Navier-Stokes equations, ergodicity has been first obtained in [40, 60, 48, 23], where the results have been further refined in [22, 12, 9, 75, 47, 41, 70, 64, 63, 5, 49]. Up to now, questions of ergodicity for stochastic Navier-Stokes equations with non-degenerate noise are an active topic of research, see [16, 30, 42, 13, 15, 44, 66].
The setup for SPDEs with locally monotone drift has been introduced by Liu and Röckner in [55]. As a unifying approach, it has sparked a lot of interest and has lead to many subsequent works, see [56, 3, 14, 36, 65, 69, 50, 1, 2, 33], among many other contributions. Examples for well-posedness covered by the approach of locally monotone drift SPDEs are the stochastic incompressible 2D Navier-Stokes equations, stochastic incompressible power law fluids, the stochastic incompressible tamed 3D Navier-Stokes equations, the stochastic 3D Leray-$\alpha$ model, as well as, quasilinear equations as the stochastic $p$-Laplace equation with non-monotone perturbation, and, with certain restrictions on the growth behavior and spatial dimension, semilinear stochastic equations as the 1D stochastic Burgers equation, and the 1D stochastic Allen-Cahn equation with double-well potential, and a large class of reaction-diffusion equations.
In this work, our main aim is to provide an abstract framework for ergodicity of locally monotone SPDEs with degenerate additive Lévy noise. Existence and uniqueness of solutions to our model have been obtained by Brzeźniak, Liu and Zhu in [14]. As our main example, we prove the weak ergodicity of the stochastic incompressible 2D Navier-Stokes equations with Dirichlet boundary conditions on a bounded domain or with periodic boundary conditions on a square without any spatial non-degeneracy condition on the noise, thus including the deterministic incompressible 2D Navier-Stokes equation. The examples of stochastic semilinear equations include the stochastic Burgers equation in 1D, however, due to our conditions, only small quadratic growth perturbations of the second order differential operator are permitted, thus ruling out the stochastic Allen-Cahn equation in 1D with double-well potential. This is not a surprise, however, as generally speaking, the locally monotone framework of Liu and Röckner [55, 56] is better suited for SPDEs where the highest order term is pseudo-monotone or strongly dissipative, as the construction of solutions is based on Faedo-Galerkin approximations and weak convergence methods, whereas semilinear drift SPDEs are often constructed via fixed point arguments, as for instance in the approach of Agresti and Veraar [3, 1, 2], which uses critical exponents and a local Lipschitz condition. Ergodicity for locally monotone drift SPDEs has been obtained in [80] for non-degenerate noise.
The methods
Our main result Theorem 2.1 below is proved in Section 5. The proof consists of the following steps. First, we derive certain a priori estimates for the solutions. These lead to the so-called $e$-property of the semigroup. To the best of our knowledge, the $e$-property for locally monotone SPDEs is proved here for the first time. It is worthwhile to note, that the $e$-property is trivial for monotone drift SPDEs with additive noise. The main tool used here is the highly useful stochastic Gronwall Lemma by Agresti and Veraar [3], which is a special case of the more general stochastic Gronwall lemma by Geiss [32]. The next step is to compare the paths of the solutions to the solution of the deterministic partial differential equation (deterministic PDE) with noise set to zero, which has an exponential or polynomial decay behavior due to our coercivity assumptions. The conditional coupling will be proved again with the help of the stochastic Gronwall Lemma [3] and a small ball property for the Lévy noise [7]. By this kind of coupling, we are able to avoid exponential martingales, which requires the existence of exponential moments. It leads to an asymptotic irreducibility property which yields the lower bound condition of Komorowski, Peszat and Szarek [46]. This procedure to prove ergodicity has been first followed by Es-Sarhir and von Renesse in [28], and has been applied to singular multi-valued monotone SPDEs by Gess and the second author in [34].
Organization of the paper
In Section 2, we present our model and the main results of this paper. In Section 3, we shall pose and discuss the main hypotheses for our results, namely Hypotheses (A)(B)(C)(D)(E)(F), and (G). Subsequently, we discuss the necessary background from Markovian semigroups and invariant measures, together with the most important auxiliary results. In Section 4, we prove Theorem 2.2, Theorem 2.3, Theorem 2.4, and Theorem 2.5. In Section 5, we prove Theorem 2.1. We collect a useful inequality in Appendix A in order to get an explicit representation of the constants appearing in the main estimates. In Appendix B, we prove a convexity result, which is used in the proof of Proposition 5.3. In Appendix C, we prove a small ball property for symmetric Lévy processes needed for the proof of Lemma 5.8.
Notation
For a metric space $\mathcal{X}$, denote the continuous and bounded real-valued functions on $\mathcal{X}$ by $C_{b}(\mathcal{X})$, equipped with the supremum norm $\|f\|_{\infty}:=\sup_{x\in\mathcal{X}}|f(x)|$. Denote the Lipschitz continuous functions from $\mathcal{X}$ to $\mathbb{R}$ by $\operatorname{Lip}(\mathcal{X})$, and denote the bounded and Lipschitz continuous functions from $\mathcal{X}$ to $\mathbb{R}$ by $\operatorname{Lip}_{b}(\mathcal{X})$. For $f\in\operatorname{Lip}(\mathcal{X})$, denote the Lipschitz constant of $f$ by $\operatorname{Lip}(f)$. $\operatorname{Lip}_{b}(\mathcal{X})$ is equipped with the norm $\|f\|_{\operatorname{Lip}_{b}}:=\operatorname{Lip}(f)+\|f\|_{\infty}$. Denote by ||$\cdot$|| ⋅ |, $\cdot$, respectively, the Euclidean norm on $\mathbb{R}^{d}$, and the Euclidean scalar product on $\mathbb{R}^{d}\times\mathbb{R}^{d}$, respectively. The transpose of a real vector or matrix is denoted by the upper index ${}^{\textup{t}}$. The adjoint of a linear operator on a Hilbert space is denoted by the upper index . For a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, we denote by $W_{0}^{1,p}({\mathcal{O}};\mathbb{R}^{k})$, $k\in\mathbb{N}$, the closure of compactly supported smooth functions $C_{0}^{\infty}({\mathcal{O}};\mathbb{R}^{k})$ in $L^{p}({\mathcal{O}};\mathbb{R}^{k})$, with respect to the Sobolev norm $\|v\|_{1,p}:=\left(\int_{\mathcal{O}}|\nabla v|^{p}\,\mathrm{d}x\right)^{1/p}$. For a Banach space $V$, denote space of linear operators from $V$ to $V$ by $L(V)$ with operator norm $\|\cdot\|_{L(V)}$. For a separable Hilbert space $H$, denote the space of Hilbert-Schmidt operators from $H$ to $H$ by $L_{2}(H)$ with Hilbert-Schmidt norm $\|\cdot\|_{2}:=\|\cdot\|_{L_{2}(H)}$. Denote by $\ell^{1}$ and $\ell^{2}$, respectively, the spaces of absolutely summable and square summable real-valued sequences, respectively. As usually, $a\wedge b$ denotes the minimum of two real numbers $a$ and $b$, and $a\vee b$ denotes the maximum of two real numbers $a$ and $b$.
2. The model and main results
We are interested in SPDEs with locally monotone drift and additive Lévy noise, which consists of a Gaussian part and an independent Poisson part with small jumps. Let $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t})_{t\geq 0},\mathbb{P})$ be a filtered probability space satisfying the standard conditions. Let $H$ be a separable Hilbert space. Let $V$ be a reflexive Banach space embedded linearly, densely, and continuously to $H$. We would like to point out that we do not assume compactness of the embedding, even if the embedding may be compact in several examples. We consider the unique strong solution $(X_{t})_{t\geq 0}$ of
(2.1) $\mathrm{d}X_{t}=A(X_{t})\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x\in H,$
where $(W_{t})_{t\geq 0}$ is a $H$-valued cylindrical Wiener process on $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t})_{t\geq 0},\mathbb{P})$, see [21]. The locally monotone nonlinear drift operator $A:V\to V^{\ast}$ and the bounded linear operator $B:H\to H$ satisfy the hypotheses in Section 3, in particular, it has finite Hilbert-Schmidt norm $\|B\|_{2}:=\|B\|_{L_{2}(H)}$. Assume that $(Z,{\mathcal{Z}},\pi)$ is a $\sigma$-finite measure space, where $\pi$ is the intensity measure of a stationary Poisson point process on $Z$. Denote by $D^{c}$ (representing small jumps) the complement of a fixed measurable set $D\in{\mathcal{Z}}$ (representing large jumps), where we assume that
(2.2) $\pi(D)<\infty.$
Denote the compensated Poisson random measure $\widetilde{N}:(0,\infty)\times\Omega\times{\mathcal{Z}}\to\mathbb{R}\cup\{-\infty,\infty\}$ by
(2.3) $\widetilde{N}((0,t]\times U)=N((0,t]\times U)-t\,\pi(U),\quad t>0,\quad U\in{\mathcal{Z}},$
where $N:[0,T]\times\Omega\times{\mathcal{Z}}\to[0,\infty]$ is a Poisson random measure with intensity $\pi$, in particular,
(2.4) $\mathbb{E}[N((0,t]\times U)]=t\,\pi(U),\quad t>0,\quad U\in{\mathcal{Z}},$
see [74, 14, 51]. Let $G:Z\to H$ be a strongly ${\mathcal{Z}}$-measurable function satisfying Hypothesis (F). For the rigorous definition of the stochastic integrals, we refer to [14, 21, 71, 11].
2.1. Main results
Our main result is the following abstract ergodicity result for SPDEs of the type (2.1) with locally monotone drift with degenerate and spatially regular Lévy noise. See Subsection 3.4 for the precise definitions of the terminology.
Theorem 2.1.
Assume that Hypotheses (A)(B)(C)(D)(E)(F) and (G) given in Section 3 hold true. Then, the stochastically continuous Markovian Feller semigroup $(P_{t})_{t\geq 0}$ associated to (2.1) satisfies the $e$-property in $H$ and is weak-mean ergodic. Moreover, it admits a unique invariant probability measure $\mu_{\ast}$ on $(H,{\mathcal{B}}(H))$ that admits finite $(\alpha+\beta)$-moments in $H$, where $\alpha\geq 2$ and $\beta\geq 0$ are as in Hypothesis (C). If $\beta=0$, $\mu_{\ast}$ admits finite $\alpha$-moments in $V$.
Under the additional assumption that $V\subset H$ is a compact embedding (which we do not assume), Hypothesis (B) already guarantees that the existence of at least one invariant measure, which can be proved by the classical method of Krylov and Bogoliubov, see [20, Theorem 3.1.1].
In Section 4, we discuss the following examples. First, for the sake of completeness, we consider the stochastic heat equation on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, with Dirichlet boundary conditions on a Lipschitz boundary,
(2.5) $\mathrm{d}X_{t}=\Delta X_{t}\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}\in L^{2}({\mathcal{O}}).$
We report the following result, which is well-known for the case of Gaussian noise, see [8, 20]. See [68] for the the case of degenerate Lévy noise.
Theorem 2.2.
Assume that Hypotheses (F) and (G) hold true. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic heat equation (2.5) with degenerate additive Lévy noise possesses a unique invariant measure $\mu_{\ast}$ with finite second moments in the stronger space $V$ in all spatial dimensions.
Furthermore, we consider the semilinear stochastic equation on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\{1,2,3\}$, $\nu>0$, with Lipschitz boundary,
(2.6) $\mathrm{d}X_{t}=\left(\nu\Delta X_{t}+\mathbf{f}(X_{t})\cdot\nabla X_{t}+g(X_{t})\right)\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}\in L^{2}({\mathcal{O}}).$
For $\mathbf{f}:=(f_{1},\ldots,f_{d})^{\textup{t}}$, where $v^{\textup{t}}$ denotes the transpose of $v$, we assume that $f_{i}\in\operatorname{Lip}_{b}(\mathbb{R})$, $i=1,\ldots,d$ for $d=1,2$. If $d=3$, we consider the case $f_{i}:=b_{i}\in L^{\infty}({\mathcal{O}})$, $i=1,\ldots,3$. Let $c_{0}>0$ be the embedding constant of the embedding $V\subset H$, see Fact 3.1 below. We assume that $g:\mathbb{R}\to\mathbb{R}$ is continuous with $g(0)=0$, and that there exist non-negative constants $s,C,K,k$ and $c\in(0,\nu c_{0}^{2})$, such that
(2.7) $|g(x)|\leq C\left(1+|x|^{2}\right),\quad x\in\mathbb{R},$
and that
(2.8) $\left(g(x)-g(y)\right)(x-y)\leq c\left(1+|y|^{s}\right)(x-y)^{2},\quad x,y\in\mathbb{R},$
such that for $d=1,2$, we assume that $s\leq 2$, and for $d=3$, we assume that $s\leq\frac{4}{3}$. For $d=1$, assume that $c<\nu c_{0}^{2}$. For $d=2,3$, assume that
(2.9) $\frac{c}{2\nu c_{0}^{2}}+\frac{\|\mathbf{f}\|_{L^{\infty}}}{\nu c_{0}}<\frac{1}{2}.$
Furthermore, we assume that
(2.10) $g(x)x\leq K+k|x|^{2},\quad x\in\mathbb{R}.$
Assume also that
(2.11) $\frac{\|\mathbf{f}\|_{L^{\infty}}}{\nu c_{0}}+\frac{k}{\nu c_{0}^{2}}<2.$
We impose Dirichlet boundary conditions. We get the following result.
Theorem 2.3.
Assume that $\mathbf{f}$ and $g$ satisfy (2.7), (2.8), (2.9), (2.10) and (2.11) and assume that Hypotheses (F) and (G) for the noise are valid. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic semilinear equations of the form (2.6) for $d\leq 3$, with degenerate additive Lévy noise possesses a unique invariant measure $\mu_{\ast}$ with finite second $V$-moments for $d=1,2$, and finite $\frac{16}{3}$th moments for $d=3$. In particular, this holds for the stochastic Burgers equation in 1D for $f_{1}(x)=x$.
For the stochastic incompressible 2D Navier-Stokes equations on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{2}$, $d\in\mathbb{N}$, with Dirichlet boundary conditions on a Lipschitz boundary $\partial{\mathcal{O}}$, with viscosity $\nu>0$,
(2.12) $\mathrm{d}X_{t}=\left(\nu\mathbf{P}\Delta X_{t}-\mathbf{P}\left[(X_{t}\cdot\nabla)X_{t}\right]\right)\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad\mathbf{P}(X_{0})\in L^{2}({\mathcal{O}};\mathbb{R}^{2}),$
where $\mathbf{P}$ denotes the Helmholtz-Leray projection on the solenoidal fields, we get the following result, which is, to our knowledge, novel for degenerate noise.
Theorem 2.4.
Assume that Hypotheses (F) and (G) for the noise hold true. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic incompressible 2D Navier-Stokes equations (2.12) with degenerate additive Lévy noise possesses a unique invariant measure $\mu_{\ast}$ with finite fourth moments.
As a generalization, one can consider the velocity field of a viscous and incompressible non-Newtonian fluid perturbed by Lévy noise with Dirichlet boundary conditions on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$ with Lipschitz boundary $\partial{\mathcal{O}}$, $d\geq 2$. Let $p>2$, $\nu>0$, and $u:{\mathcal{O}}\to\mathbb{R}^{d}$. Define
$e(u):{\mathcal{O}}\to\mathbb{R}^{d}\otimes\mathbb{R}^{d},\quad e_{i,j}(u):=\frac{\partial_{i}u_{j}+\partial_{j}u_{i}}{2},\quad 1\leq i,j\leq d,$
and
$\tau(u):{\mathcal{O}}\to\mathbb{R}^{d}\otimes\mathbb{R}^{d},\quad\tau(u):=2\nu(1+|e(u)|)^{p-2}e(u),$
where $\otimes$ denotes the usual tensor product. Consider the stochastic power-law fluid equations
(2.13) $\mathrm{d}X_{t}=\left(\nu\mathbf{P}(\operatorname{div}(\tau(X_{t})))-\mathbf{P}\left[(X_{t}\cdot\nabla)X_{t}\right]\right)\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad\mathbf{P}(X_{0})\in L^{2}({\mathcal{O}};\mathbb{R}^{d}).$
We obtain the following result.
Theorem 2.5.
Assume $p\geq\frac{2+d}{2}\vee 2$, and that Hypotheses (F) and (G) hold true. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic shear thickening power-law fluid equations (2.13) with degenerate additive Lévy noise possesses a unique invariant measure $\mu_{\ast}$ with finite $\frac{2p}{2p-d}$th moments.
We note that our results remain true for periodic boundary conditions, that is, when one replaces ${\mathcal{O}}$ by the flat torus $\mathbb{T}^{d}$ in all of the examples above, see Remark 4.3.
3. Preliminaries and hypotheses
The following hypotheses are modifications and partial extensions, needed for our ergodicity result, of the hypotheses for the variational well-posedness result for SPDEs with locally monotone drift in finite time from [55, 56] for Gaussian noise, which has been extended to Lévy noise in [14]. See also [36, 65, 3, 50, 69, 1, 2] for further extensions, which cover examples that are out of the scope of this paper; for instance the 3D stochastic tamed Navier-Stokes equation, the stochastic $p$-Laplace equation, the stochastic Allen-Cahn equation, and the stochastic Cahn-Hilliard equation, where the drift of the latter SPDE satisfies bounds which have been called fully locally monotone in [69].
Let $H$ be a separable Hilbert space. Let $V$ be a reflexive Banach space embedded linearly, densely, and continuously to $H$. Note that we identify $H$ with its Hilbert space dual $H^{\ast}$ by the Riesz isometry, so that we obtain a Gelfand triple
$V\subset H\equiv H^{\ast}\subset V^{\ast},$
where $V^{\ast}$ denotes the topological dual of $V$. We shall use the notation $\langle u,v\rangle$ both for $u\in V$ and $v\in V^{\ast}$, where it denotes the evaluation of a dual element, and for $u,v\in H$, where it denotes the Hilbert space inner product of $u$ and $v$ in $H$, where, in particular, for any $u\in V$ is considered as an element of $H$ by the continuous embedding, and in the latter case both meanings of the notation coincide.
Fact 3.1 (Embedding).
As the embedding $V\subset H$ is linear and bounded, there exists a constant $c_{0}>0$ such that
$\|x\|_{V}\geq c_{0}\|x\|_{H}\quad\textrm{ for any }\quad x\in V.$
3.1. Hypotheses on the drift operator
Consider the following set of hypotheses.
Hypothesis (A) (Hemicontinuity).
For every $x,y,z\in V$ the map
$\mathbb{R}\ni\lambda\mapsto\langle A(x+\lambda y),z\rangle\quad\textrm{ is continuous.}$.
Hypothesis (B) (Coercivity).
There exist constants $\delta_{1}>0$, and $\alpha\geq 2$, such that
$2\langle A(x),x\rangle\leq-\delta_{1}\|x\|^{\alpha}_{V}\quad\textrm{ for all }\quad x\in V.$
Hypothesis (C) (Local monotonicity).
There exist constants $\delta_{2}>0$, $C_{2}\geq 0$, and $\beta\geq 0$ such that
$2\langle A(x)-A(y),x-y\rangle\leq[-\delta_{2}+\rho(y)]\|x-y\|^{2}_{H}\quad\textrm{ for all }\quad x,y\in V,$
where $\alpha\geq 2$ is given in Hypothesis (B) and $\rho:V\to[0,\infty)$ is a measurable function, locally bounded in $V$, satisfying
$0\leq\rho(y)\leq C_{2}\|y\|^{\alpha}_{V}\|y\|^{\beta}_{H}\quad\textrm{ for any}\quad y\in V.$
Hypothesis (D) (Growth).
There exists a constant $K>0$ satisfying
(3.1) $\|A(x)\|_{V^{\ast}}^{\frac{\alpha}{\alpha-1}}\leq K(1+\|x\|_{V}^{\alpha})(1+\|x\|_{H}^{\beta})\quad\textrm{ for all }\quad x\in V,$
where $\alpha$ and $\beta$ are given in Hypotheses (B) and (C), respectively.
Hypothesis (E) (Cone condition).
There exist constants $\delta_{4}>0$ and $C_{4}\in\mathbb{R}$ such that
(3.2) $2\langle A(x),x\rangle\leq C_{4}-\delta_{4}\|A(x)\|_{V^{\ast}}\quad\textrm{ for any }\quad x\in V.$
Note that Hypothesis (E) is implied by
$2\langle A(x),x\rangle\leq C_{4}-\delta_{4}\|A(x)\|_{V^{\ast}}^{q}\quad\textrm{ for any }\quad x\in V$
for some $q\geq 1$.
Remark 3.2.
Note that above, especially Hypothesis (C) and Hypothesis (E) differ from the standard assumptions for locally monotone drift SPDEs [56]. To obtain global coupling at infinity (the so-called $e$-property, see Definition 3.7 below), we need to obtain a priori estimates independent of the terminal time $T>0$. To get this, we need to assume $\delta_{2}>0$, and we cannot permit constant positive perturbations of the drift. Essentially, this means that our drift has a strongly dissipative part. Hypothesis (E) is a cone condition that is of technical nature, but is easy to verify for many examples where the growth of the lower order nonlinearities can be controlled by the potential of the leading order term, see the examples in Section 4 below. Hypothesis (E) can for instance be verified with the help of the conditions of Lemma 4.1 and Lemma 4.2 below. Note that we need Hypothesis (E) only in the proof of Lemma 5.8 below.
3.2. Hypotheses on the noise
Hypothesis (F) ($H$-regularity of the noise).
The noise coefficient $B:H\to H$ is linear, symmetric, bounded, and positive definite, and satisfies
(3.3) $B\in L_{2}(H),$
where $\|B\|_{2}:=\|B\|_{L_{2}(H)}$. The intensity measure $\pi$ of $N$ satisfies (2.2). The strongly measurable jump coefficient $G:Z\to H$ satisfies
$\mathfrak{g}_{0}:=\int_{D^{c}}\|G(z)\|_{H}^{2}\,\pi(\mathrm{d}z)<\infty,$
and
$\mathfrak{g}_{\beta}:=\int_{D^{c}}\|G(z)\|_{H}^{\beta+2}\,\pi(\mathrm{d}z)<\infty,$
where $\beta\geq 0$ is as in Hypothesis (C).
Hypothesis (G) ($V$-Regularity of the noise).
For any $T>0$, the driving process of the noise
(3.4) $L_{t}:=B\,W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\{t\},\mathrm{d}z),\quad t\in[0,T],$
where $(W_{t})_{t\in[0,T]}$ and $\left(\widetilde{N}(\{t\}\times U)\right)_{t\in[0,T],\;U\in\mathcal{Z}}$ are independent, has a representation
(3.5) $L_{t}=\sum_{k=1}^{\infty}\sqrt{\lambda_{k}}e_{k}l_{t}^{k},\quad t\in[0,T],$
where $e_{k}\in V$, $k\in\mathbb{N}$ are elements of a complete orthonormal system of $H$, $(l^{k}_{t})_{t\in[0,T]}$ are independent and identically distributed (i.i.d.) one-dimensional symmetric Lévy processes with finite second moments, such that $l^{1}_{1}$ has characteristic triplet $(1,\pi_{0},0)$, and $\lambda_{k}\geq 0$ with $\{\lambda_{k}\}\in\ell^{1}$, and where the series converges weakly in $L^{2}(\Omega,{\mathcal{F}},\mathbb{P};H)$ for every $t\in[0,T]$.
Moreover, we assume that there exist $\sigma_{k}\geq c_{0}^{2}>0$, $k\in\mathbb{N}$, with
(3.6) $\|e_{k}\|_{V}\leq\sqrt{\sigma_{k}},\quad k\in\mathbb{N},$
such that
$\{\sigma_{k}\lambda_{k}\}\in\ell^{1}$
and
[…]
Remark 3.3.
  1. (i)
    Hypothesis (G) is needed to verify a small ball probability for the noise in $V$, see Appendix C, which is only used in the coupling estimate Lemma 5.8 below. When $V$ is a separable Hilbert space, Parseval’s identity yields that for verifying Hypothesis (G) it is sufficient to assume that $\{\sqrt{\sigma_{k}}e_{k}\}$ is a complete orthonormal system of $V$. In this case $\|e_{k}\|_{V}=\sqrt{\sigma_{k}}$ for $k\in\mathbb{N}$. See [6, Theorem 4.7] for cylindrical representations of Lévy processes of the above type.
  2. (ii)
    Note that $\{\lambda_{k}\}\in\ell^{1}$ implies $\{\sqrt{\lambda_{k}}\}\in\ell^{2}$. Furthermore, if $G\equiv 0$, by Hypothesis (G), the operator $B:H\to H$, defined by
    $Bv:=\sum_{k=1}^{\infty}\sqrt{\lambda_{k}}\langle v,e_{k}\rangle e_{k},$
    has finite Hilbert-Schmidt norm
    $\|B\|_{L_{2}(H)}=\left(\sum_{k=1}^{\infty}\|Be_{k}\|^{2}_{H}\right)^{1/2}=\|\{\sqrt{\lambda_{k}}\}\|_{\ell^{2}}<\infty.$
    Hence, in this case, Hypothesis (G) implies Hypothesis (F).
3.3. Existence and uniqueness of solutions
We start recalling the definition of solution for (2.1) that we shall use here.
Definition 3.4.
An $H$-valued càdlàg [^1] That is, right-continuous with left limits. $({\mathcal{F}}_{t})_{t\in[0,T]}$-adapted process $(X_{t})_{t\in[0,T]}$ is called a solution to (2.1) if for its $\mathrm{d}t\otimes\mathbb{P}$-equivalence class $\widehat{X}$, we have
  1. (i)
    $\widehat{X}\in L^{\alpha}([0,T];V)\cap L^{2}([0,T];H)$, $\mathbb{P}$-a.s.
  2. (ii)
    For any $V$-valued $({\mathcal{F}}_{t})_{t\in[0,T]}$-progressively measurable $\mathrm{d}t\otimes\mathbb{P}$-version $\overline{X}$ of $\widehat{X}$ it holds that
[…]
Now, we recall the existence and uniqueness result given in [14], and [50], respectively. Compare also [3] for a novel alternative approach to locally monotone SPDEs using interpolation spaces.
Proposition 3.5 (Existence and uniqueness).
Let $T>0$ be fixed. Assume that Hypotheses (A), (B), (C), (D) and (F) hold true. Then for any $x\in L^{\beta+2}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$, there exists a unique solution to (2.1) in the sense of Definition 3.4.
Proof.
By our hypotheses, the conditions of [14, Theorem 1.2] are satisfied. ∎
3.4. Invariant measures
Recall that the Markovian semigroup $(P_{t})_{t\geq 0}$ associated to (2.1) acts as follows
$P_{t}F(x):=\mathbb{E}[F(X_{t}^{x})]\quad\textrm{ for any }\quad F\in B_{b}(H)\quad\textrm{ and }\quad x\in H,$
where $B_{b}(H):=\{F:H\to\mathbb{R}:\,F\textrm{ is bounded and Borel measurable}\}$. See [56, Proposition 4.3.5] for a proof of the Markov property in the Gaussian noise case. The Lévy noise case follows by taking essentially the same steps. See also [34, Section 6.4] for a discussion of the Markov property for additive noise SPDEs. For a semigroup $(P_{t})_{t\geq 0}$, we define the dual semigroup $(P_{t}^{\ast})_{t\geq 0}$ acting on ${\mathcal{M}}_{1}(H,{\mathcal{B}}(H)):=\{\mu:{\mathcal{B}}(H)\to[0,1]:\,\mu\textrm{ is a probability measure}\}$ by
$P_{t}^{\ast}\mu(A):=\int_{H}P_{t}1_{A}(x)\,\mu(\mathrm{d}x)\quad\textrm{ for any }\quad A\in{\mathcal{B}}(H),$
where ${\mathcal{B}}(H)$ denote the Borel sets of $H$ and $1_{A}$ denotes the indicator function of the set $A$, see [20] for details.
Definition 3.6 (Stochastically continuous Feller semigroup).
We say that $(P_{t})_{t\geq 0}$ is a stochastically continuous Feller semigroup if for every $x\in H$ and every $r>0$ it follows that
$\lim_{t\to 0}P_{t}^{\ast}\delta_{x}(B(x,r))=1\quad\textrm{ and }\quad P_{t}(C_{b}(H))\subset C_{b}(H),$
where $B(x,r):=\{y\in H:\|y\|_{H}<r\}$, $\delta_{x}$ denotes the Dirac delta measure at $x$ and $C_{b}(H):=\{F:H\to\mathbb{R}:\,F\textrm{ is continuous and bounded}\}$.
Definition 3.7 ($e$-property).
We say that the semigroup $(P_{t})_{t\geq 0}$ satisfies the $e$-property if for every $F\in\operatorname{Lip}_{b}(H)$, for every $x\in H$, and every $\varepsilon>0$, there exists $\delta>0$ such that
$|P_{t}F(x)-P_{t}F(y)|<\varepsilon\quad\textrm{ for all }\quad y\in B(x,\delta)\quad\textrm{ and }\quad t\geq 0,$
where $\operatorname{Lip}_{b}(H):=\{F:H\to\mathbb{R}:\,F\textrm{ is Lipschitz and bounded}\}$.
Definition 3.8 (Invariant measure).
A measure $\mu_{\ast}\in{\mathcal{M}}_{1}(H,{\mathcal{B}}(H))$ is said to be invariant for the semigroup $(P_{t})_{t\geq 0}$ if $P^{\ast}_{t}\mu_{\ast}=\mu_{\ast}$ for all $t\geq 0$.
We recall the following concepts defined e.g. in [46].
Definition 3.9 (Weak-mean ergodicity, weak law of large numbers).
We say that $\{P_{t}\}_{t\geq 0}$ is weak-mean ergodic if there exists a Borel probability measure $\mu_{\ast}$ on ${\mathcal{B}}(H)$, such that
[…]
where the limit is in the sense of weak convergence of probability measures.
We say that the weak law of large numbers holds for $\{P_{t}\}_{t\geq 0}$, for a function $F\in\operatorname{Lip}_{b}(H)$ and for a probability measure $\mu$ on ${\mathcal{B}}(H)$ if
$\mathbb{P}_{\mu}\text{-}\lim_{T\to+\infty}\frac{1}{T}\int_{0}^{T}F(X_{t})\,\mathrm{d}t=\int_{H}F\,\mathrm{d}\mu_{\ast},$
where $\mu_{\ast}$ denotes the invariant measure of $\{P_{t}\}_{t\geq 0}$ and $\{X_{t}\}_{t\geq 0}$ denotes the Markov process related to $\{P_{t}\}_{t\geq 0}$ whose initial distribution is $\mu$ and whose path measure is $\mathbb{P}_{\mu}$, and where the convergence takes place in $\mathbb{P}_{\mu}$-probability.
As noted in [46, Remark 3], (3.8) implies uniqueness of the invariant measure. Let us recall [46, Theorem 2], which is an extension of the so-called lower bound technique by Lasota and Szarek [52]. Define
$Q_{T}\mu:=\frac{1}{T}\int_{0}^{T}P^{\ast}_{s}\mu\,\mathrm{d}s$
and write $Q_{T}(x,\cdot):=Q_{T}\delta_{x}$.
Theorem 3.10 (Komorowski–Peszat–Szarek).
Assume that $(P_{t})_{t\geq 0}$ has the $e$-property and that there exists $z\in H$ such that for every bounded set $J$ and every $\delta>0$, we have
(3.9) $\inf_{x\in J}\liminf_{T\to\infty}Q_{T}(x,B(z,\delta))>0.$
Suppose further that for every $\varepsilon>0$ and every $x\in H$, there exists a bounded Borel set $K\subset H$ such that
(3.10) $\liminf_{T\to\infty}Q_{T}(x,K)>1-\varepsilon.$
Then there exists a unique invariant measure $\mu_{\ast}$ for $(P_{t})_{t\geq 0}$ such that the semigroup $(P_{t})_{t\geq 0}$ is weak-mean ergodic and the weak law of large numbers holds.
Proof.
See [46, Theorem 2]. ∎
4. Examples
Let us start with two lemmas discussing Hypothesis (E). If $A$ is of subgradient type on $V$, we may obtain Hypothesis (E) more easily as follows. We refer to [24] for the terminology of the subgradient of a convex functional.
Lemma 4.1.
Suppose that $A=-\partial\Phi$, that is, $A$ is equal to the negative subgradient of a lower semi-continuous convex functional $\Phi:V\to\mathbb{R}$ such that $\inf_{u\in V}\Phi(u)>-\infty$. Then Hypothesis (E) is satisfied.
Proof.
As $V$ is reflexive, the proof of [34, Proposition 7.1] can be adapted verbatim. ∎
For operators with a subgradient principal part, Hypothesis (E) can be checked as follows.
Lemma 4.2.
If $A=A_{0}+F$ where $A_{0}=-\partial\Phi$ is the subgradient of a lower semi-continuous convex functional $\Phi:V\to\mathbb{R}$ such that $\inf_{u\in V}\Phi(u)>-\infty$, and $F:V\to V^{\ast}$ is a strongly measurable (nonlinear) operator. Suppose that for some constants $\kappa_{1}\in\mathbb{R}\setminus\{0\}$, $K_{1}\in\mathbb{R}$ that for every $u\in V$,
(4.1) $\|F(u)\|_{V^{\ast}}\leq\kappa_{1}\langle A_{0}u,u\rangle+K_{1}.$
Furthermore, suppose that for some constants $K_{2},\kappa_{2}\in\mathbb{R}$, for every $u\in V$,
(4.2) $2\langle F(u),u\rangle\leq\kappa_{2}\|F(u)\|_{V^{\ast}}+K_{2},$
or that
(4.3) $2\langle F(u),u\rangle\leq\kappa_{2}\langle A_{0}u,u\rangle+K_{2},$
where, in the second case, we additionally assume that $\kappa_{2}\in(-2,\infty)$, whenever $\kappa_{1}<0$. Then $A$ satisfies Hypothesis (E).
Proof.
By Lemma 4.1, there exists $\tilde{\delta}_{4}>0$ and $\tilde{C}_{4}\in\mathbb{R}$ with
$2\langle A_{0}u,u\rangle\leq\tilde{C}_{4}-\tilde{\delta}_{4}\|A_{0}u\|_{V^{\ast}}$
for every $u\in V$. If assumption (4.1) and (4.2) hold, and if $\kappa_{1}>0$, for $R>\kappa_{2}\vee 0$,
$2\langle Au,u\rangle=$ $2\langle A_{0}u,u\rangle+2\langle F(u),u\rangle$
$\leq$ $\left(2+R\kappa_{1}\right)\langle A_{0}u,u\rangle-R\kappa_{1}\langle A_{0}u,u\rangle+\kappa_{2}\|F(u)\|_{V^{\ast}}+K_{2}$
$\leq$ $\left(1+\frac{R\kappa_{1}}{2}\right)\tilde{C}_{4}+RK_{1}+K_{2}-\left(1+\frac{R\kappa_{1}}{2}\right)\tilde{\delta}_{4}\|A_{0}u\|_{V^{\ast}}-(R-\kappa_{2})\|F(u)\|_{V^{\ast}}$
$\leq$ $\left(1+\frac{R\kappa_{1}}{2}\right)\tilde{C}_{4}+RK_{1}+K_{2}-\left(\left(1+\frac{R\kappa_{1}}{2}\right)\tilde{\delta}_{4}\wedge(R-\kappa_{2})\right)\|A_{0}u+F(u)\|_{V^{\ast}}.$
Hence Hypothesis (E) holds with $\delta_{4}=\left(1+\frac{R\kappa_{1}}{2}\right)\tilde{\delta}_{4}\wedge(R-\kappa_{2})$ and $C_{4}=\left(1+\frac{R\kappa_{1}}{2}\right)\tilde{C}_{4}+RK_{1}+K_{2}$.
On the other hand, if $\kappa_{1}<0$,
$2\langle Au,u\rangle=$ $2\langle A_{0}u,u\rangle+2\langle F(u),u\rangle$
$\leq$ $\langle A_{0}u,u\rangle+\langle A_{0}u,u\rangle+(\kappa_{2}\vee 0)\|F(u)\|_{V^{\ast}}+K_{2}$
$\leq$ $\langle A_{0}u,u\rangle+(1+(\kappa_{2}\vee 0))\langle A_{0}u,u\rangle+K_{1}+K_{2}$
$\leq$ $\frac{\tilde{C}_{4}}{2}+K_{1}+K_{2}-\frac{\tilde{\delta}_{4}}{2}\|A_{0}u\|_{V^{\ast}}-\frac{1+(\kappa_{2}\vee 0)}{|\kappa_{1}|}\|F(u)\|_{V^{\ast}}+(K_{1}\vee 0)\frac{1+(\kappa_{2}\vee 0)}{|\kappa_{1}|}$
$\leq$ $\frac{\tilde{C}_{4}}{2}+K_{2}+(K_{1}\vee 0)\frac{1+(\kappa_{2}\vee 0)+|\kappa_{1}|}{|\kappa_{1}|}-\left(\frac{\tilde{\delta}_{4}}{2}\wedge\frac{1+(\kappa_{2}\vee 0)}{|\kappa_{1}|}\right)\|A_{0}u+F(u)\|_{V^{\ast}}.$
Hence Hypothesis (E) holds with $\delta_{4}=\frac{\tilde{\delta}_{4}}{2}\wedge\frac{1+(\kappa_{2}\vee 0)}{|\kappa_{1}|}$ and $C_{4}=\frac{\tilde{C}_{4}}{2}+K_{2}+(K_{1}\vee 0)\frac{1+(\kappa_{2}\vee 0)+|\kappa_{1}|}{|\kappa_{1}|}$.
If assumptions (4.1) and (4.3) hold, if $\kappa_{1}>0$, for $R>\frac{|\kappa_{2}|}{\kappa_{1}}$,
$2\langle Au,u\rangle=$ $2\langle A_{0}u,u\rangle+2\langle F(u),u\rangle$
$\leq$ $\left(2+R\kappa_{1}+\kappa_{2}\right)\langle A_{0}u,u\rangle-R\kappa_{1}\langle A_{0}u,u\rangle+K_{2}$
$\leq$ $\left(1+\frac{R\kappa_{1}+\kappa_{2}}{2}\right)\tilde{C}_{4}+RK_{1}+K_{2}-\left(1+\frac{R\kappa_{1}+\kappa_{2}}{2}\right)\tilde{\delta}_{4}\|A_{0}u\|_{V^{\ast}}-R\|F(u)\|_{V^{\ast}}$
$\leq$ $\left(1+\frac{R\kappa_{1}+\kappa_{2}}{2}\right)\tilde{C}_{4}+RK_{1}+K_{2}-\left(\left(1+\frac{R\kappa_{1}+\kappa_{2}}{2}\right)\tilde{\delta}_{4}\wedge R\right)\|A_{0}u+F(u)\|_{V^{\ast}}.$
Hence Hypothesis (E) holds with $\delta_{4}=\left(1+\frac{R\kappa_{1}+\kappa_{2}}{2}\right)\tilde{\delta}_{4}\wedge R$ and $C_{4}=\left(1+\frac{\kappa_{1}+\kappa_{2}}{2}\right)\tilde{C}_{4}+RK_{1}+K_{2}$.
On the other hand, if $\kappa_{1}<0$, $\kappa_{2}\in(-2,0)$, for $\varepsilon\in(0,2+\kappa_{2})$,
$2\langle Au,u\rangle=$ $2\langle A_{0}u,u\rangle+2\langle F(u),u\rangle$
$\leq$ $(2+\kappa_{2}-\varepsilon)\langle A_{0}u,u\rangle+\varepsilon\langle A_{0}u,u\rangle+K_{2}$
$\leq$ $\frac{2+\kappa_{2}-\varepsilon}{2}\tilde{C}_{4}-\frac{\tilde{\delta}_{4}(2+\kappa_{2}-\varepsilon)}{2}\|A_{0}u\|_{V^{\ast}}-\frac{\varepsilon}{|\kappa_{1}|}\|F(u)\|_{V^{\ast}}+(K_{1}\vee 0)\frac{\varepsilon}{|\kappa_{1}|}+K_{2}$
$\leq$ $\frac{2+\kappa_{2}-\varepsilon}{2}\tilde{C}_{4}+(K_{1}\vee 0)\frac{\varepsilon}{|\kappa_{1}|}+K_{2}-\left(\frac{\tilde{\delta}_{4}(2+\kappa_{2}-\varepsilon)}{2}\wedge\frac{\varepsilon}{|\kappa_{1}|}\right)\|A_{0}u+F(u)\|_{V^{\ast}}.$
Hence Hypothesis (E) holds with $\delta_{4}=\frac{\tilde{\delta}_{4}(2+\kappa_{2}-\varepsilon)}{2}\wedge\frac{\varepsilon}{|\kappa_{1}|}$ and $C_{4}=\frac{2+\kappa_{2}-\varepsilon}{2}\tilde{C}_{4}+(K_{1}\vee 0)\frac{\varepsilon}{|\kappa_{1}|}+K_{2}$. ∎
Remark 4.3.
We note that the results of this section also hold in an analog manner if one replaces the bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$ with Dirichlet boundary conditions with the flat torus $\mathbb{T}^{d}:=\mathbb{R}^{d}/(2\pi\mathbb{Z}^{d})$, that is, periodic boundary conditions, combined with the requirement that $\int_{\mathbb{T}^{d}}X_{t}\,\mathrm{d}x=0$, $\mathbb{P}$-a.e. for every $t\geq 0$.
4.1. Stochastic heat equation
Let us prove Theorem 2.2. Consider the stochastic heat equation on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$ with Lipschitz boundary,
(4.4) $\mathrm{d}X_{t}=\Delta X_{t}\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x,$
where $x\in H:=L^{2}({\mathcal{O}})$. We impose Dirichlet boundary conditions, so that $V:=W^{1,2}_{0}({\mathcal{O}})$. Due to Poincaré’s inequality, we shall equip $V$ with the norm
$\|v\|_{V}:=\left(\int_{\mathcal{O}}|\nabla v|^{2}\,\mathrm{d}x\right)^{1/2},\quad v\in V.$
Note that the embedding constant $c_{0}$ is the inverse of the Poincaré constant of ${\mathcal{O}}$.
Hypothesis (A):
The hemicontinuity has been proved in [67, Example 4.1.7].
Hypothesis (B):
The coercivity has been proved in [67, Example 4.1.7], with $\alpha=2$ and $\delta_{1}=2$.
Hypothesis (C):
The monotonicity has been proved in [67, Example 4.1.7] with $\alpha=2$, $\beta=0$, $\delta_{2}=2$, $\rho\equiv 0$, and $C_{2}=0$.
Hypothesis (D):
The growth condition has been proved in [67, Example 4.1.7].
Hypothesis (E):
Hypothesis (E) follows from Lemma 4.1 for the functional $\Phi=\frac{1}{2}\|\cdot\|_{V}^{2}$.
Hence, Theorem 2.2 follows immediately from Theorem 2.1.
4.2. Semilinear stochastic equations
Let us prove Theorem 2.3. Consider the stochastic semilinear equation on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, $\nu>0$, with Lipschitz boundary,
(4.5) $\mathrm{d}X_{t}=\nu\Delta X_{t}\,\mathrm{d}t+\sum_{i=1}^{d}f_{i}(X_{t})\partial_{i}X_{t}\,\mathrm{d}t+g(X_{t})\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x,$
where $x\in H:=L^{2}({\mathcal{O}})$. We impose Dirichlet boundary conditions, so that $V:=W^{1,2}_{0}({\mathcal{O}})$. Due to Poincaré’s inequality, we shall equip $V$ with the norm
$\|v\|_{V}:=\left(\int_{\mathcal{O}}|\nabla v|^{2}\,\mathrm{d}x\right)^{1/2},\quad v\in V.$
Note that the embedding constant $c_{0}$ is the inverse of the Poincaré constant of ${\mathcal{O}}$. We assume that $f_{i}\in\operatorname{Lip}_{b}(\mathbb{R})$, $i=1,\ldots,d$ for $d=1,2$. For $d=3$, we consider the case that $f_{i}:=b_{i}\in L^{\infty}({\mathcal{O}})$, $i=1,\ldots,3$. Denote $\mathbf{f}:=(f_{1},\ldots,f_{d})^{\textup{t}}$. We assume that $g:\mathbb{R}\to\mathbb{R}$ is continuous with $g(0)=0$, and that there exist $s,C,c\geq 0$, such that
(4.6) $|g(x)|\leq C\left(1+|x|^{2}\right),\quad x\in\mathbb{R},$
and
(4.7) $\left(g(x)-g(y)\right)(x-y)\leq c\left(1+|y|^{s}\right)(x-y)^{2},\quad x,y\in\mathbb{R}.$
Assume for $d=3$ that $s=\frac{4}{3}$. For $d=1$, assume that $c<\nu c_{0}^{2}$. For $d=2,3$, assume that
(4.8) $\frac{c}{2\nu c_{0}^{2}}+\frac{\|\mathbf{f}\|_{L^{\infty}}}{\nu c_{0}}<\frac{1}{2}.$
Let us also assume that there exist $K,k\geq 0$ with
(4.9) $g(x)x\leq K+k|x|^{2},\quad x\in\mathbb{R}.$
Assume that
(4.10) $\frac{\|\mathbf{f}\|_{L^{\infty}}}{\nu c_{0}}+\frac{k}{\nu c_{0}^{2}}<2.$
Note that assumption (4.9) and (4.10) are needed to prove Hypothesis (E).
Hypothesis (A):
The hemicontinuity can be proved as in [56, Lemma 5.1.6 and Example 5.1.7].
Hypothesis (B):
The coercivity can be proved as in [56, Lemma 5.1.6 and Example 5.1.7] with $\alpha=2$.
Hypothesis (C):
The monotonicity for $\alpha=2$ can be proved as in [56, Lemma 5.1.6 and Example 5.1.7]. As we have to give some attention to the constant $\delta_{2}>0$, we will discuss part of the arguments here. Let $d=1$ and $c<\nu c_{0}^{2}$. Then, by [56, Equation (5.15) and Example 5.1.8],
$\begin{split}&2\langle A(u)-A(v),u-v\rangle\\ \leq&-2\nu\|u-v\|_{V}^{2}+2\operatorname{Lip}(\mathbf{f})\left(\|u-v\|_{V}\|v\|_{L^{\infty}}\|u-v\|_{H}+\|v\|_{V}\|u-v\|_{L^{4}}^{2}\right)\\ &+2c\left(1+\|v\|_{L^{\infty}}^{s}\right)\|u-v\|_{H}^{2}.\end{split}$
Now, due to
(4.11) $\|u\|_{L^{4}}^{2}\leq 2\|u\|_{L^{2}}\|\nabla u\|_{L^{2}}$
for every $u\in W_{0}^{1,2}({\mathcal{O}})$ in $d=1,2$, we get that for any $\varepsilon\in(0,2\nu c_{0}^{2}-2c)$, there exists a constant $C=C(d,{\mathcal{O}},c_{0},\operatorname{Lip}(\mathbf{f}),\varepsilon,c,s)>0$ with
$2\langle A(u)-A(v),u-v\rangle\leq-(2\nu c_{0}^{2}-2c-\varepsilon)\|u-v\|_{H}^{2}+C\|v\|_{V}^{2}\|u-v\|_{H}^{2}.$
Let $d=2$. Then, by [56, Equations (5.15) and (5.16), and Example 5.1.8],
$\begin{split}&2\langle A(u)-A(v),u-v\rangle\\ \leq&-2\nu\|u-v\|_{V}^{2}+4\|\mathbf{f}\|_{\infty}\|u-v\|_{V}\|u-v\|_{H}+2\operatorname{Lip}(\mathbf{f})\|v\|_{V}\|u-v\|_{L^{4}}^{2}\\ &+2c\left(1+\|v\|_{L^{2s}}^{s}\right)\|u-v\|_{H}^{2}\\ \leq&-\left(2\nu c_{0}^{2}-4c_{0}\|\mathbf{f}\|_{\infty}\right)\|u-v\|_{H}^{2}+2\operatorname{Lip}(\mathbf{f})\|v\|_{V}\|u-v\|_{L^{4}}^{2}\\ &+2c\left(1+\|v\|_{L^{2s}}^{s}\right)\|u-v\|_{H}^{2}.\end{split}$
Now, by (4.8) and by (4.11), we get that for any $\varepsilon\in(0,2\nu c_{0}^{2}-2c-4c_{0}\|\mathbf{f}\|_{\infty})$, there exists a constant $C=C(d,{\mathcal{O}},c_{0},\operatorname{Lip}(\mathbf{f}),\varepsilon,c,s)>0$ with
$\begin{split}&2\langle A(u)-A(v),u-v\rangle\\ \leq&-\left(2\nu c_{0}^{2}-4c_{0}\|\mathbf{f}\|_{\infty}-\varepsilon\right)\|u-v\|_{H}^{2}+C\|v\|_{V}^{2}\|u-v\|_{H}^{2}.\end{split}$
Hence, in $d=1,2$, Hypothesis (C) holds with $\beta=0$ and $\alpha=2$.
Let $d=3$. Then by [56, Lemma 5.1.6 and Example 5.1.8], and taking into account the interpolation inequality
$\|u\|_{L^{4}}^{2}\leq 2\sqrt{2}\|u\|_{H}^{\frac{1}{2}}\|u\|_{V}^{\frac{3}{2}},\quad u\in W_{0}^{1,2}({\mathcal{O}}),$
this case follows by similar arguments, with $\alpha=2$ and $\beta=\frac{10}{3}$, see [56, Example 5.1.8 (3)].
Hypothesis (D):
Hypothesis (D) follows as in [56, Lemma 5.1.6 and Example 5.1.7].
Hypothesis (E):
Note that $\nu\Delta u=-\partial\Phi(u)$ for the functional $\Phi=\frac{\nu}{2}\|\cdot\|_{V}^{2}$. As $\mathbf{f}$ is essentially bounded, we get that
$\langle(\mathbf{f}(u)\cdot\nabla u),u\rangle\leq\|\mathbf{f}\|_{L^{\infty}}\|u\|_{V}\|u\|_{H}\leq\frac{\|\mathbf{f}\|_{L^{\infty}}}{c_{0}}\|u\|_{V}^{2}=-\frac{\|\mathbf{f}\|_{L^{\infty}}}{\nu c_{0}}\langle\nu\Delta u,u\rangle.$
Furthermore,
$\|\mathbf{f}(u)\cdot\nabla u\|_{V^{\ast}}\leq\frac{1}{c_{0}}\|\mathbf{f}(u)\cdot\nabla u\|_{H}$
$\leq$ $\frac{\|\mathbf{f}\|_{L^{\infty}}}{c_{0}}\|u\|_{V}\leq\frac{\|\mathbf{f}\|_{L^{\infty}}}{c_{0}}\|u\|_{V}^{2}+\frac{\|\mathbf{f}\|_{L^{\infty}}}{c_{0}}$
$=$ $-\frac{\|\mathbf{f}\|_{L^{\infty}}}{\nu c_{0}}\langle\nu\Delta u,u\rangle+\frac{\|\mathbf{f}\|_{L^{\infty}}}{c_{0}}.$
By (4.9), we have that
$\langle g(u),u\rangle\leq k\|u\|^{2}_{L^{2}}+K\leq\frac{k}{c_{0}^{2}}\|u\|^{2}_{V}+K.$
Hence,
$\langle g(u),u\rangle\leq-\frac{k}{\nu c_{0}^{2}}\langle\nu\Delta u,u\rangle+K.$
Similarly, by the Sobolev embedding theorem, there exists $c_{4}>0$ with
$\|g(u)\|_{V^{\ast}}\leq\frac{1}{c_{0}}\|g(u)\|_{L^{2}}\leq\frac{C}{c_{0}}\|u\|_{L^{4}}^{2}+C\leq\frac{c_{4}C}{c_{0}}\|u\|_{V}^{2}+C=-\frac{c_{4}C}{\nu c_{0}}\langle\nu\Delta u,u\rangle+C.$
Now by (4.10), Hypothesis (E) follows now from Lemma 4.2.
Consequently, Theorem 2.3 follows directly from Theorem 2.1.
4.3. Stochastic incompressible 2D Navier-Stokes equations
Let us prove Theorem 2.4. The deterministic incompressible 2D Navier-Stokes equation can be formulated as
$\partial_{t}u(t)$ $=\nu\Delta u(t)-(u(t)\cdot\nabla)u(t)-\nabla p(t)+f(t),$
$\nabla\cdot u(t)$ $=0,$
where $u:[0,T]\times{\mathcal{O}}\to\mathbb{R}^{2}$ denotes the velocity field of an incompressible Newtonian fluid with viscosity $\nu>0$, $p:[0,T]\times{\mathcal{O}}\to\mathbb{R}$ denotes the pressure, and $f:[0,T]\times{\mathcal{O}}\to\mathbb{R}^{2}$ denotes an external force, where ${\mathcal{O}}\subset\mathbb{R}^{2}$ is a bounded domain with sufficiently smooth boundary, see [77]. We employ Dirichlet boundary conditions. Define
$V:=\{v\in W^{1,2}_{0}({\mathcal{O}};\mathbb{R}^{2})\;\colon\;\nabla\cdot v=0\text{ a.e. in }{\mathcal{O}}\}.$
Due to Poincaré’s inequality, we shall equip $V$ with the norm
$\|v\|_{V}:=\left(\int_{\mathcal{O}}|\nabla v|^{2}\,\mathrm{d}x\right)^{1/2},\quad v\in V.$
We define $H$ as the closure of $V$ in $L^{2}({\mathcal{O}};\mathbb{R}^{2})$ w.r.t. the standard $L^{2}$-norm. $H$ is a closed Hilbert subspace of $L^{2}({\mathcal{O}};\mathbb{R}^{2})$, so the Helmholtz-Leray projection
$\mathbf{P}:L^{2}({\mathcal{O}};\mathbb{R}^{2})\to H$
is given by the orthogonal projection. The Stokes operator with viscosity constant $\nu$ is given by
$A_{0}:W^{2,2}({\mathcal{O}};\mathbb{R}^{2})\cap V\to H,\quad A_{0}u=\nu\mathbf{P}\Delta u.$
Denote $F(u):=F(u,u)$, where
$F(u,v)=-\mathbf{P}[(u\cdot\nabla)v],\quad u,v\in V.$
Denote the extensions $A_{0}:V\to V^{\ast}$ and $F:V\times V\to V^{\ast}$ by the same respective symbols. Note that these extensions can be defined by duality for Lipschitz boundary. Consider the stochastic incompressible 2D Navier-Stokes equation with viscosity $\nu>0$ driven by degenerate additive Lévy noise
(4.12) $\mathrm{d}X_{t}=(A_{0}X_{t}+F(X_{t}))\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x,$
where $B\in L_{2}(H)$ and $D\in{\mathcal{Z}}$, and $(Z,{\mathcal{Z}},\pi)$ is a measure space as in the hypotheses. Existence and uniqueness of solutions to (2.12) have been discussed in [14, Example 2.6]. Let us verify our Hypotheses in the Gelfand triple
$V\subset H\equiv H^{\ast}\subset V^{\ast}.$
It is well-known, see [77] or [61, Lemma 2.1] that
$\|u\|_{L^{4}({\mathcal{O}},\mathbb{R}^{2})}^{4}\leq 2\|u\|_{H}^{2}\|u\|_{V}^{2},\quad u\in V.$
Note that the embedding constant $c_{0}$ is the inverse of the Poincaré constant of ${\mathcal{O}}$.
Hypothesis (A):
The hemicontinuity has been proved in [14, Example 2.6].
Hypothesis (B):
The coercivity has been proved [14, Example 2.6], with $\alpha=2$ and $\delta_{1}=2\nu$.
Hypothesis (C):
By [61, Lemma 2.2], noting that $\langle F(u),u\rangle=0$ for $u\in V$, we get for all $u,v\in V$,
$2\langle F(u)-F(v),u-v\rangle$
$=$ $-2\langle F(u,u-v),v\rangle+2\langle F(v,u-v),v\rangle$
$=$ $-2\langle F(u-v),v\rangle$
$\leq$ $4\|u-v\|_{V}^{3/2}\|u-v\|^{1/2}_{H}\|v\|_{L^{4}({\mathcal{O}},\mathbb{R}^{2})}^{4}$
$\leq$ $\nu\|u-v\|_{V}^{2}+\frac{64}{\nu^{3}}\|v\|_{L^{4}({\mathcal{O}},\mathbb{R}^{2})}^{4}\|u-v\|_{H}^{2}$
as in [14, Equation (2.20)] for
$\rho(v):=\frac{64}{\nu^{3}}\|v\|_{L^{4}({\mathcal{O}};\mathbb{R}^{2})}^{4}\leq\frac{128}{\nu^{3}}\|v\|_{H}^{2}\|u\|_{V}^{2}.$
Hence
$2\langle A_{0}u+F(u)-A_{0}v-F(v),u-v\rangle\leq-\nu\|u-v\|^{2}_{V}+\rho(v)\|u-v\|_{H}^{2}.$
We get that Hypothesis (B) holds with $\alpha=\beta=2$, $\delta_{2}=\nu c_{0}^{2}$ and $\rho(v)=\frac{64}{\nu^{3}}\|v\|_{L^{4}({\mathcal{O}},\mathbb{R}^{2})}^{4}$ and $C_{2}=\frac{128}{\nu^{3}}$.
Hypothesis (D):
Hypothesis (D) holds by [14, Example 2.6].
Hypothesis (E):
Note that $A_{0}$ has the convex potential $\frac{1}{2}\|A_{0}^{1/2}u\|_{H}^{2}$, $u\in V$. Furthermore, we have that $\langle F(u),u\rangle=0$ for every $u\in V$. By Hölder’s inequality,
$|\langle F(u),v\rangle|\leq\|u\|_{L^{4}({\mathcal{O}};\mathbb{R}^{2})}^{2}\|v\|_{V},\quad u,v\in V,$
which yields by the Sobolev embedding $V\subset L^{4}({\mathcal{O}};\mathbb{R}^{2})$ for some constant $C>0$ that
$\|F(u)\|_{V^{\ast}}\leq C\|u\|_{V}^{2}=\frac{C}{\nu}\langle A_{0}^{1/2}u,A_{0}^{1/2}u\rangle=-\frac{C}{\nu}\langle A_{0}u,u\rangle,$
for every $u\in V$. Hypothesis (E) follows now from Lemma 4.2.
Then, Theorem 2.4 follows directly from Theorem 2.1.
4.4. Stochastic shear thickening incompressible power-law fluids
Let us prove Theorem 2.5. Consider the velocity field of a viscous and incompressible non-Newtonian fluid perturbed by Lévy noise with Dirichlet boundary conditions on a sufficiently smooth bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, with outward unit normal $n$ on $\partial{\mathcal{O}}$, $d\geq 2$. Let $p>2$, $\nu>0$, and assume that $p\geq\frac{2+d}{2}$. The case $p>2$ is called shear thickening. For $u:{\mathcal{O}}\to\mathbb{R}^{d}$, define
$e(u):{\mathcal{O}}\to\mathbb{R}^{d}\otimes\mathbb{R}^{d},\quad e_{i,j}(u):=\frac{\partial_{i}u_{j}+\partial_{j}u_{i}}{2},\quad 1\leq i,j\leq d,$
and
$\tau(u):{\mathcal{O}}\to\mathbb{R}^{d}\otimes\mathbb{R}^{d},\quad\tau(u):=2\nu(1+|e(u)|)^{p-2}e(u).$
Let
$V:=\{u\in W_{0}^{1,p}({\mathcal{O}};\mathbb{R}^{d})\;\colon\;\operatorname{div}(u)=0\text{ a.e. in }{\mathcal{O}}\},$
and
$H:=\{u\in L^{2}({\mathcal{O}};\mathbb{R}^{d})\;\colon\;\operatorname{div}(u)=0\text{ a.e. in }{\mathcal{O}},\;u\cdot n=0\text{ on }\partial{\mathcal{O}}\}.$
Denote the Helmholtz-Leray projection by $\mathbf{P}:L^{2}({\mathcal{O}};\mathbb{R}^{d})\to H$. Define the nonlinear $p$-Stokes operator by
$A_{0}:W^{2,p}({\mathcal{O}};\mathbb{R}^{d})\cap V\to H,\quad A_{0}(u):=\mathbf{P}(\operatorname{div}(\tau(u))),$
and define the convection term $F:(W^{2,p}({\mathcal{O}};\mathbb{R}^{d})\cap V)\times(W^{2,p}({\mathcal{O}};\mathbb{R}^{d})\cap V)\to H$ as before in the case of the Navier-Stokes equation as
$F(u,v):=-\mathbf{P}[(u\cdot\nabla)v],\quad F(u):=F(u,u).$
Denote the extensions $A_{0}:V\to V^{\ast}$ and $F:V\times V\to V^{\ast}$ by the same respective symbols. Note that these extensions can be defined by duality for Lipschitz boundary. Consider the stochastic power-law fluid equations
(4.13) $\mathrm{d}X_{t}=(A_{0}X_{t}+F(X_{t}))\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x.$
Existence and uniqueness of this equation has been discussed in [14, Example 2.9].
Hypothesis (A):
The hemicontinuity has been proved in [14, Example 2.9].
Hypothesis (B):
The coercivity has been proved [14, Example 2.9], with $\alpha=p$ for some $\delta_{1}>0$.
Hypothesis (C):
By [14, Example 2.9], we get for all $u,v\in V$, that there exists $C>0$, such that for any $\varepsilon>0$, there exists $C(\varepsilon)>0$ with
$2\langle A_{0}u+F(u)-A_{0}v-F(v),u-v\rangle\leq-(C-\varepsilon)\|u-v\|^{2}_{V}+\rho(v)\|u-v\|_{H}^{2},$
where $\rho(v):=C_{\varepsilon}\|v\|_{V}^{\frac{2p}{2p-d}}$, and $\rho(v)\leq C_{\varepsilon}\|v\|_{V}^{p}\|v\|_{H}^{\beta}$ for $\beta=\frac{2p}{2p-d}-p=\frac{p(2-2p+d)}{2p-d}$.
Hypothesis (D):
Hypothesis (D) holds by [14, Example 2.9].
Hypothesis (E):
Note that $A_{0}$ is the subgradient of the continuous convex [^2] Note that convexity follows from the fact that $f(x):=2\nu(p^{-1}(1+|x|)^{p}-(p-1)^{-1}(1+|x|)^{p-1})$ has the non-negative second derivative $f^{\prime\prime}(x)=2\nu(|x|+1)^{p-3}(1+(p-1)|x|)$. potential on $V$
$v\mapsto 2\nu\int_{{\mathcal{O}}}\left(\frac{1}{p}(1+|e(v)|)^{p}-\frac{1}{p-1}(1+|e(v)|)^{p-1}\right)\,\mathrm{d}x,\quad v\in V.$
By the Riesz-Thorin interpolation inequality and the Sobolev embedding theorem, for $q=\frac{dp}{d-p}$ and $\gamma=\frac{d}{(d+2)p-2d}$, for some constant $C>0$,
$\|v\|_{L^{\frac{2p}{p-1}}}\leq\|v\|^{\gamma}_{L^{q}}\|v\|^{1-\gamma}_{H}\leq C\|v\|_{V},\quad v\in V.$
Note that by Korn’s inequality [58, Theorem 1.10 (p. 196)] and the previous inequality, for some $C>0$, which might change from line to line,
(4.14) $\begin{split}-\langle A_{0}v,v\rangle=&C+C\int_{{\mathcal{O}}}|e(v)|^{p}\,\mathrm{d}x\\ \geq&C+C\|v\|_{V}^{p}\\ \geq&C+C\|v\|^{p}_{L^{\frac{2p}{p-1}}}\\ \geq&C+C\|v\|^{2}_{L^{\frac{2p}{p-1}}},\quad v\in V,\end{split}$
as $p>2$. Also,
(4.15) $\|F(v)\|_{V^{\ast}}\leq\|v\|^{2}_{L^{\frac{2p}{p-1}}},\quad v\in V,$
see [14, Example 2.9]. Therefore, Hypothesis (E) is satisfied by Lemma 4.2, noting that $\langle F(u),u\rangle=0$ for $u\in V$.
Now, Theorem 2.5 follows from Theorem 2.1.
5. Proof of Theorem 2.1
In this section, we shall verify Theorem 2.1 by proving that the conditions for Theorem 3.10 hold.
5.1. Main estimates
We start with defining some constants.
Definition 5.1.
Assume that Hypotheses (A)(B)(C)(D), and (F) are valid. For any $\lambda_{i}\in(0,1)$, $i=1,2,3,4$, and $\lambda_{0}\in(0,1)$ with $\sum_{i=0}^{4}\lambda_{i}=1$, let
(5.1) $\begin{split}&c_{1}:=\frac{\beta+2}{2}\left(\frac{\beta}{(\alpha+\beta)\lambda_{1}c_{0}^{\alpha}\delta_{1}}\right)^{\frac{\beta}{\alpha}}\frac{\alpha}{\alpha+\beta}\left(\left(1+18(\beta+2)\right)\|B\|^{2}_{2}+2\beta\|B\|_{L(H)}^{2}\right)^{\frac{\alpha+\beta}{\alpha}},\\ &c_{2}:=\left(\frac{2(\alpha-2)}{(\beta+2)(\alpha+\beta)\lambda_{2}c_{0}^{\alpha}\delta_{1}}\right)^{\frac{2+\beta}{\alpha-2}}\frac{\alpha-2}{\alpha+\beta}(\gamma\|B\|_{2}^{2})^{\frac{\alpha+\beta}{\alpha-2}},\\ &c_{3}:=\frac{\beta+2}{2}\left(\frac{\beta}{(\alpha+\beta)\lambda_{1}c_{0}^{\alpha}\delta_{1}}\right)^{\frac{\beta}{\alpha}}\frac{\alpha}{\alpha+\beta}(18(\beta+2)\mathfrak{g}_{0})^{\frac{\alpha+\beta}{\alpha}},\\ &c_{4}:=\left(\frac{2\beta}{(\beta+2)(\alpha+\beta)\lambda_{4}c_{0}^{\alpha}\delta_{1}}\right)^{\frac{\beta}{\alpha}}\frac{\alpha}{\alpha+\beta}(C_{\beta}\mathfrak{g}_{0})^{\frac{\alpha+\beta}{\alpha}},\\ &c_{5}:=C_{\beta}\mathfrak{g}_{\beta},\end{split}$
if $\alpha>2$, where
$C_{\beta}:=((2^{\beta-3}\vee 1)\beta(\beta+2))\vee((2^{\beta-1}\vee 1)(\beta+2)),\quad\beta>0,$
and $C_{0}:=\frac{1}{2}$ are as in Lemma A.1. Define $c_{1},c_{2},c_{4}$, and $c_{5}$ similarly if $\alpha=2$, and $c_{2}:=0$. Set
$\mathfrak{c}:=\sum_{i=1}^{5}c_{i}.$
Remark 5.2.
Note that $\mathfrak{c}=0$ if and only if $B\equiv G\equiv 0$, that is, we are considering the deterministic equation without noise. In this case, in the proof of Proposition 5.3 below, we can choose $\lambda_{0}=1$ and $\lambda_{i}=0$, $i=1,2,3,4$. In the pure Gaussian case, that is, $G\equiv 0$ and $B\not=0$, we may set $\lambda_{3}=\lambda_{4}=c_{3}=c_{4}=c_{5}=0$. Analogously, in the pure jump case $B\equiv 0$ and $G\not=0$, we may set $\lambda_{1}=\lambda_{2}=c_{1}=c_{2}=0$.
Let us prove the main a priori estimate for solutions to (2.1), given our hypotheses.
Proposition 5.3 (Main estimate).
Assume that Hypotheses (A)(B)(C)(D), and (F) are valid. Let $T>0$ be fixed. Let $\alpha\geq 2,\beta\geq 0$ be as in Hypothesis (C). Then for any initial datum $x\in L^{\beta+2}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$, and every $0\leq s\leq t\leq T$, the solution $(X_{t})_{t\in[0,T]}$ to (2.1) satisfies
(5.2) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t}\|X_{r}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t}e^{\gamma(r-t)}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|X_{s}\|_{H}^{\beta+2}e^{\gamma(s-t)}\right]+\frac{\mathfrak{c}}{\gamma}\left(1-e^{\gamma(s-t)}\right),\end{split}$
for some $\gamma\in(0,\delta_{2})$, where $\delta_{1}>0$ is as in Hypothesis (B), and $\delta_{2}>0$ is as in Hypothesis (C), and where $\lambda_{0}\in(0,1]$ and $\mathfrak{c}\geq 0$ are as in Definition 5.1, where $\lambda_{0}=1$ and $\mathfrak{c}=0$ if and only if $B\equiv G\equiv 0$. In particular, $\mathfrak{c}$ does not depend on $T$.
Proof.
In order to rigorously apply Itô’s formula, we construct approximating solutions to (2.1) via projections to finite dimensional subspaces using the so-called Faedo-Galerkin’s method.
In the sequel, we define the notation of the aforementioned projections. Let $\{e_{1},e_{2},\ldots,\}\subset V$ be an orthonormal basis for $H$ such that $\operatorname{span}\{e_{1},e_{2},\ldots,\}$ is dense in $V$. For each $n\in\mathbb{N}$, we denote $H_{n}:=\operatorname{span}\{e_{1},\ldots,e_{n}\}$ and set $Q_{n}:V^{\ast}\to H_{n}$ be the linear projection operator defined by
(5.3) $Q_{n}y:=\sum_{j=1}^{n}\langle y,e_{j}\rangle e_{j}\quad\textrm{ for any }\quad y\in V^{\ast}.$
We note that $Q_{n}|_{H}$ is the orthogonal projection from $H$ onto $H_{n}$, which, by abuse of notation, is denoted by the same symbol.
We then denote for each $t\geq 0$
$W^{(n)}_{t}:=\sum_{j=1}^{n}\langle W_{t},e_{j}\rangle e_{j}=Q_{n}W_{t}.$
Bearing all this in mind, for a fixed $x\in L^{\beta+2}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$ and for each $n\in\mathbb{N}$, we consider the solution $(X^{(n),x_{n}}_{t})_{t\geq 0}$ of the following stochastic differential equation on the finite dimensional space $H_{n}$:
(5.4) $\begin{split}\mathrm{d}X^{(n),x_{n}}_{t}&=Q_{n}A(X_{t}^{(n),x_{n}})\,\mathrm{d}t+Q_{n}B\,\mathrm{d}W_{t}+\int_{D^{c}}Q_{n}G(w)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}w),\quad t\geq 0,\\ X_{0}^{(n),x_{n}}&=Q_{n}x=:x_{n}.\end{split}$
We point out that there exists a unique probabilistically strong solution of (5.4), see for instance [4, 37].
We remark that $X^{(n),x_{n}}_{t}\in H_{n}\subset V$ for all $t\geq 0$. For all $n\in\mathbb{N}$ we note that $\|Q_{n}y\|_{H_{n}}\leq\|y\|_{H}$ for all $y\in H$, and $\|z\|_{H_{n}}=\|z\|_{H}$ for all $z\in H_{n}$. Since we are interested in moments estimates, we introduce the following localization argument. Let $T>0$ be a fixed time-horizon and take $R>\|x_{n}\|_{H_{n}}+1$. We then define the stopping time
(5.5) $\tau^{(n),x_{n}}_{R}:=\inf\left\{s\geq 0:\|X^{(n),x_{n}}_{s}\|_{V}>R\right\}\wedge T.$
By (5.5) we have $\|X^{(n),x_{n}}_{t}\|_{V}\leq R$ for $0\leq t<\tau^{(n),x_{n}}_{R}$ and hence by Fact 3.1 we obtain $\|X^{(n),x_{n}}_{t}\|_{H}\leq c^{-1}_{0}R$ for $0\leq t<\tau^{(n),x_{n}}_{R}$. Similarly as in the proof of [14, Lemma 4.2] or in the proof of [69, Theorem 2.8], we have
(5.6) $\tau^{(n),x_{n}}_{R}\uparrow T\quad\mathbb{P}\textrm{-a.s. as $R\uparrow\infty$}$
and
(5.7) $\lim\limits_{R\to\infty}\sup_{n\in\mathbb{N}}\mathbb{P}(\tau^{(n),x_{n}}_{R}<T)=0.$
For the sake of readability, let us abbreviate $\tau_{R}:=\tau^{(n),x_{n}}_{R}$, and $X^{(n)}:=X^{(n),x_{n}}$. By Itô’s formula [38, 62, 39], or [74, Section 2.8], see also [14, proof of Lemma 4.2], it follows that $\mathbb{P}$-a.s.
(5.8) $\begin{split}\|X_{t}^{(n)}\|^{\beta+2}_{H}e^{\gamma t}=&\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma s}+\frac{\beta+2}{2}\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\left(2\langle A(X_{r}^{(n)}),X_{r}^{(n)}\rangle\right)\,\mathrm{d}r\\ &+\frac{\beta+2}{2}\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\|Q_{n}BQ_{n}\|_{2}^{2}\,\mathrm{d}r\\ &+\beta(\beta+2)\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta-2}_{H}\|(Q_{n}BQ_{n})^{\ast}X^{(n),x_{n}}_{r}\|^{2}_{H}\,\mathrm{d}r\\ &+(\beta+2)\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle\\ &+(\beta+2)\int_{s}^{t}\int_{D^{c}}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X^{(n),x_{n}}_{r},Q_{n}G(w)\rangle\,\widetilde{N}(\mathrm{d}r,\mathrm{d}w)\\ &+\int_{s}^{t}\int_{D^{c}}e^{\gamma r}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|^{\beta+2}_{H}-\|X_{r}^{(n)}\|^{\beta+2}_{H}\right.\\ &\qquad\left.-(\beta+2)\|X_{r}^{(n)}\|_{H}^{\beta}\langle X_{r}^{(n)},Q_{n}G(w)\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)\\ &+\gamma\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r\end{split}$
for all $0\leq s\leq t\leq T$, where $\widetilde{N}$ denotes the compensated Poisson random measure defined in (2.3). And thus, for any $0\leq s\leq t\leq T$, $\mathbb{P}$-a.s., by Hypothesis (B), noting that $B$ is symmetric,
$\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}+\delta_{1}\frac{\beta+2}{2}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r$
$\leq$ $\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-(t\wedge\tau_{R}))}$
$+\frac{\beta+2}{2}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|_{2}^{2}\,\mathrm{d}r$
$+(\beta+2)e^{-\gamma(t\wedge\tau_{R})}\sup_{s\leq u\leq t\wedge\tau_{R}}\left|\int_{s}^{u}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle\right|$
$+(\beta+2)e^{-\gamma(t\wedge\tau_{R})}\sup_{s\leq u\leq t\wedge\tau_{R}}\left|\int_{s}^{u}\int_{D^{c}}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}G(w)\rangle\,\widetilde{N}(\mathrm{d}r,\mathrm{d}w)\right|$
$+\beta(\beta+2)\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|^{2}_{L(H)}\,\mathrm{d}r$
$+\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.$
$\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)$
$+\gamma\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r,$
where the penultimate term is non-negative by the subgradient inequality (see e.g. [24]) and the fact that $u\mapsto\|u\|^{\beta+2}_{H}$ is convex with Fréchet differential $u\mapsto(\beta+2)\|u\|^{\beta}_{H}\langle u,h\rangle$, $h\in H$. By the Burkholder-Davis-Gundy inequality [45, Theorem 1], and by Young’s inequality,
$\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]+\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r$
$\leq$ $\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-(t\wedge\tau_{R}))}\right]$
$+\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|_{2}^{2}\,\mathrm{d}r$
$+3(\beta+2)\mathbb{E}\left[\left(\int_{s}^{t\wedge\tau_{R}}e^{2\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|B\|_{2}^{2}\,\mathrm{d}r\right)^{\frac{1}{2}}\right]$
$+3(\beta+2)\mathbb{E}\left[\left(\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{2\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|G(w)\|_{H}^{2}\,\pi(\mathrm{d}w)\,\mathrm{d}r\right)^{\frac{1}{2}}\right]$
$+\beta(\beta+2)\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|^{2}_{L(H)}\,\mathrm{d}r$
$+\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.$
$\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)$
$+\gamma\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r$
$\leq$ $\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-(t\wedge\tau_{R}))}\right]$
$+\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|_{2}^{2}\,\mathrm{d}r$
$+9(\beta+2)^{2}\mathbb{E}\left[\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|_{2}^{2}\,\mathrm{d}r\right]+\frac{1}{4}\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]$
$+9(\beta+2)^{2}\mathbb{E}\left[\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|G(w)\|_{H}^{2}\,\pi(\mathrm{d}w)\,\mathrm{d}r\right]$
$+\frac{1}{4}\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]$
$+\beta(\beta+2)\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|^{2}_{L(H)}\,\mathrm{d}r$
$+\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.$
$\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)$
$+\gamma\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r,$
Hence, for $\lambda_{0}\in(0,1)$ as in Definition 5.1, and by Fubini’s theorem,
$\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]$
$+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r+(1-\lambda_{0})I_{t\wedge\tau_{R},s}^{(n)}$
$\leq$ $\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-(t\wedge\tau_{R}))}\right]$
$+\left(\left(\frac{1}{2}(\beta+2)+9(\beta+2)^{2}\right)\|B\|_{2}^{2}+\beta(\beta+2)\|B\|_{L(H)}^{2}\right)\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\,\mathrm{d}r$
$+9(\beta+2)^{2}\mathfrak{g}_{0}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\,\mathrm{d}r$
$+\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.$
$\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)$
$+\gamma\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r,$
where
$I_{t,s}^{(n)}:$ $=c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}e^{-\gamma t}\mathbb{E}\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\alpha+\beta}_{H}\,\mathrm{d}r,$
for any $0\leq s\leq t\leq T$ and any $n\in\mathbb{N}$.
Our hypotheses guarantee that all constants in Definition 5.1 are finite. Recall Young’s inequality for products, that is, for any $p,q\in(1,\infty)$ satisfying $p^{-1}+q^{-1}=1$ it follows that
(5.9) $|xy|\leq\varepsilon|x|^{p}+\frac{(p\varepsilon)^{1-q}}{q}|y|^{q}\quad\text{ for any }\quad x,y\in\mathbb{R}\quad\textrm{ and }\quad\varepsilon>0.$
If $\alpha>2$, we get that for any $x\in H$, and any $\lambda_{i}\in[0,1)$, $i=1,2,3,4$, and $\lambda_{0}\in(0,1]$ with $\sum_{i=0}^{4}\lambda_{i}=1$, $c_{i}\geq 0$, $i=1,\ldots,5$ as in Definition 5.1, with
(5.10) $\begin{split}&\left(\left(\frac{\beta+2}{2}+9(\beta+2)^{2}\right)\|B\|_{2}^{2}+\beta(\beta+2)\|B\|_{L(H)}^{2}\right)\|x\|^{\beta}_{H}\leq\lambda_{1}c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}\|x\|^{\alpha+\beta}_{H}+c_{1},\\ &\gamma\|x\|^{\beta+2}_{H}\|B\|_{2}^{2}\leq\lambda_{2}c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}\|x\|^{\alpha+\beta}_{H}+c_{2},\\ &9(\beta+2)^{2}\mathfrak{g}_{0}\|x\|^{\beta}_{H}\leq\lambda_{1}c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}\|x\|^{\alpha+\beta}_{H}+c_{3},\\ &C_{\beta}\mathfrak{g}_{0}\|x\|^{\beta}_{H}\leq\lambda_{4}c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}\|x\|^{\alpha+\beta}_{H}+c_{4},\\ &C_{\beta}\mathfrak{g}_{\beta}=c_{5},\end{split}$
for any $\gamma>0$. If $\alpha=2$, the inequalities for $c_{1},c_{3}$, and $c_{4}$ follow from (5.9), and $c_{2}=0$ follows by choosing $\gamma\in(0,\delta_{2})$ such that
(5.11) $\gamma\|B\|_{2}^{2}\leq\lambda_{2}c_{0}^{2}\delta_{1}\frac{\beta+2}{2}.$
As a consequence, we have that for $c_{i}\geq 0$, $i=1,2,3$, as in Definition 5.1, that $\mathbb{P}$-a.s.
(5.12) $\begin{split}&\left(\left(\frac{\beta+2}{2}+9(\beta+2)^{2}\right)\|B\|_{2}^{2}+\beta(\beta+2)\|B\|_{L(H)}^{2}\right)\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\,\mathrm{d}r\\ \leq&\lambda_{1}I_{{t\wedge\tau_{R}},s}^{(n)}+c_{1}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r,\\ &\gamma\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r\leq\lambda_{2}I_{{t\wedge\tau_{R}},s}^{(n)}+c_{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r,\\ &9(\beta+2)^{2}\mathfrak{g}_{0}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\,\mathrm{d}r\leq\lambda_{3}I_{{t\wedge\tau_{R}},s}^{(n)}+c_{3}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r,\end{split}$
We claim that $\mathbb{P}$-a.s.
(5.13) $\begin{split}&\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.\\ &\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,\pi(\mathrm{d}w)\,\mathrm{d}r\\ \leq&\lambda_{4}I_{{t\wedge\tau_{R}},s}^{(n)}+(c_{4}+c_{5})\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r.\end{split}$
By Lemma A.1, we have for $C_{\beta}=((2^{\beta-3}\vee 1)\beta(\beta+2))\vee((2^{\beta-1}\vee 1)(\beta+2))$, $\beta>0$, and $C_{0}=\frac{1}{2}$ that
$\|x+w\|_{H}^{\beta+2}-\|x\|^{\beta+2}_{H}-(\beta+2)\|x\|^{\beta}_{H}\left\langle x,w\right\rangle\leq C_{\beta}\left(\|x\|_{H}^{\beta}\|w\|_{H}^{2}+\|w\|^{\beta+2}_{H}\right).$
Hence, by Fubini’s theorem,
(5.14) $\begin{split}&\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.\\ &\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,\pi(\mathrm{d}w)\,\mathrm{d}r\\ \leq&C_{\beta}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}\|_{H}^{\beta}\|Q_{n}G(w)\|_{H}^{2}+\|Q_{n}G(w)\|^{\beta+2}_{H}\right]\,\pi(\mathrm{d}w)\,\mathrm{d}r\\ \leq&C_{\beta}\mathfrak{g}_{0}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|_{H}^{\beta}\,\mathrm{d}r+C_{\beta}\mathfrak{g}_{\beta}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r.\end{split}$
The claim follows now by (5.10). Thus, noting that, by definition, $1-\lambda_{0}=\sum_{i=1}^{4}\lambda_{i}$, and recalling that $\mathfrak{c}:=\sum_{i=1}^{5}c_{i}$,
(5.15) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-{t\wedge\tau_{R}})}\right]+\mathfrak{c}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r\\ =&\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-{t\wedge\tau_{R}})}\right]+\frac{\mathfrak{c}}{\gamma}\left(1-\mathbb{E}\left[e^{\gamma(s-(t\wedge\tau_{R}))}\right]\right).\end{split}$
By (5.6), (5.7), and Fatou’s Lemma, we get that
(5.16) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t}e^{\gamma(r-t)}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-t)}\right]+\frac{\mathfrak{c}}{\gamma}\left(1-e^{\gamma(s-t)}\right).\end{split}$
By [14, Lemma 4.4], there exists a subsequence $\{n_{k}\}$ such that $Q_{n_{k}}X^{(n_{k}),x_{n_{k}}}\rightharpoonup^{\ast}X^{x}$ weakly in $L^{\infty}([0,T];L^{2}(\Omega;H))$ and weakly in $L^{\alpha}([0,T]\times\Omega,{\mathcal{B}}{\mathcal{F}},\mathrm{d}t\otimes\mathbb{P};V)$ as $k\to\infty$, where
:=${\mathcal{B}}{\mathcal{F}}:=\{A\subset[0,T]\times\Omega\,\colon\,\forall t\in[0,T],A\cap([0,t]\times\Omega)\in{\mathcal{B}}([0,t])\otimes{\mathcal{F}}_{t}\}$
is the $\sigma$-field of progressively measurable sets on $[0,T]\times\Omega$. By Lemma B.2, combined with Fatou’s lemma, we can pass to the limit $k\to\infty$ and get
(5.17) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t}\|X_{r}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t}e^{\gamma(r-t)}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|X_{s}\|_{H}^{\beta+2}e^{\gamma(s-t)}\right]+\frac{\mathfrak{c}}{\gamma}\left(1-e^{\gamma(s-t)}\right).\end{split}$
for $\mathrm{d}t$-almost every $0\leq s\leq t\leq T$. The proof is complete. As the right-hand side depends continuously on $s$ and $t$, we get the result for every $0\leq s\leq t\leq T$ by a simple approximation argument. ∎
By investigating the above proof, we get, by similar arguments, the following:
Corollary 5.4.
Assume that Hypotheses (A)(B)(C)(D)(E), and (F) hold true. Let $T>0$ be fixed. Let $\alpha\geq 2,\beta\geq 0$ be as in Hypothesis (C). Then for any initial datum $X_{0}\in L^{\beta+2}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$, and every $0\leq s\leq t\leq T$, the solution $(X_{t})_{t\in[0,T]}$ to (2.1) satisfies
(5.18) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t}\|X_{r}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\leq\mathbb{E}\left[\|X_{s}\|_{H}^{\beta+2}\right]+\mathfrak{c}(t-s),\end{split}$
where $\delta_{1}>0$ is as in Hypothesis (B), and $\delta_{2}>0$ is as in Hypothesis (C), and where $\lambda_{0}\in(0,1]$ and $\mathfrak{c}\geq 0$ are as in Definition 5.1, where $\lambda_{0}=1$ and $\mathfrak{c}=0$ if and only if $B\equiv G\equiv 0$. In particular, $X\in L^{\infty}([0,T];L^{\beta+2}(\Omega,{\mathcal{F}},\mathbb{P};H))$, and $\mathfrak{c}$ does not depend on $T$.
5.2. Properties of the semigroup
We collect some basic properties of the semigroup. Let us verify the $e$-property of the semigroup $(P_{t})_{t\geq 0}$ first.
Proposition 5.5 ($e$-property).
Assume that Hypotheses (A)(B)(C)(D)(E)(F), and (G) are valid. Let $\varepsilon>0$, $F\in\operatorname{Lip}_{b}(H)$, $x\in H$. Then there exists $\delta>0$ such that for any $y\in H$ with $\|x-y\|<\delta$, and any $t\geq 0$, we have that
$|P_{t}F(x)-P_{t}F(y)|<\varepsilon.$
Proof.
Note that if $C_{2}=0$, the $e$-property follows as in [34, Equation (5.4)]. Let $\varepsilon>0$, $x,y\in H$, $\|x-y\|<\delta$, $F\in\operatorname{Lip}_{b}(H)$, $t\geq 0$. Let $(X^{x}_{t})_{t\geq 0}$, $(X^{y}_{t})_{t\geq 0}$ denote the solutions to (2.1) with $X_{0}^{x}=x$, and $X^{y}_{0}=y$, respectively. W.l.o.g. $\|F\|_{\infty}>0$ and $\|F\|_{\operatorname{Lip}}>0$. We claim that
(5.19) $|P_{t}F(x)-P_{t}F(y)|<\frac{\varepsilon}{3}+2\|F\|_{\infty}\mathbb{P}\left(\left\{\|X_{t}^{x}-X_{t}^{y}\|^{2}_{H}\geq\frac{\varepsilon^{2}}{9\|F\|_{\operatorname{Lip}}^{2}}\right\}\right).$
Let $J_{\varepsilon}:=\{|F(X_{t}^{x})-F(X_{t}^{y})|<\frac{\varepsilon}{3}\}$. Then,
(5.20) $\begin{split}&|P_{t}F(x)-P_{t}F(y)|\\ \leq&\mathbb{E}[|F(X_{t}^{x})-F(X_{t}^{y})|]\\ =&\mathbb{E}[|F(X_{t}^{x})-F(X_{t}^{y})|1_{J_{\varepsilon}}]+\mathbb{E}[|F(X_{t}^{x})-F(X_{t}^{y})|1_{J_{\varepsilon}^{c}}]\\ \leq&\frac{\varepsilon}{3}+\mathbb{E}[|F(X_{t}^{x})-F(X_{t}^{y})|1_{J_{\varepsilon}^{c}}]\\ \leq&\frac{\varepsilon}{3}+2\|F\|_{\infty}\mathbb{P}(J_{\varepsilon}^{c})\\ =&\frac{\varepsilon}{3}+2\|F\|_{\infty}\mathbb{P}\left(\left\{|F(X_{t}^{x})-F(X_{t}^{y})|\geq\frac{\varepsilon}{3}\right\}\right)\\ \leq&\frac{\varepsilon}{3}+2\|F\|_{\infty}\mathbb{P}\left(\left\{\|X_{t}^{x}-X_{t}^{y}\|^{2}_{H}\geq\frac{\varepsilon^{2}}{9\|F\|_{\operatorname{Lip}}^{2}}\right\}\right),\end{split}$
which proves (5.19). Let us investigate the tail probability. By Itô’s formula and Hypothesis (C), we have for $0\leq s\leq t$ and $\gamma\in(0,\delta_{2})$ as in (5.10) that
$\|X_{t}^{x}-X_{t}^{y}\|_{H}^{2}e^{\gamma t}$
$\leq$ $\|X_{s}^{x}-X_{s}^{y}\|_{H}^{2}e^{\gamma s}+2\int_{s}^{t}e^{\gamma r}\langle A(X_{r}^{x})-A(X_{r}^{y}),X_{r}^{x}-X_{r}^{y}\rangle\,\mathrm{d}r+\gamma\int_{s}^{t}e^{\gamma r}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\,\mathrm{d}r$
$\leq$ $\|X_{s}^{x}-X_{s}^{y}\|_{H}^{2}e^{\gamma s}+C_{2}\int_{s}^{t}e^{\gamma r}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\|X_{r}^{x}\|_{V}^{\alpha}\|X_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r+(\gamma-\delta_{2})\int_{s}^{t}e^{\gamma r}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\,\mathrm{d}r.$
For any pair of stopping times $0\leq\lambda\leq\Lambda\leq t$, we get that $\mathbb{P}$-a.s.
$\sup_{\lambda\leq\tau\leq\Lambda}\|X_{\tau}^{x}-X_{\tau}^{y}\|_{H}^{2}+(\delta_{2}-\gamma)\int_{\lambda}^{\Lambda}e^{\gamma(r-\Lambda)}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\,\mathrm{d}r$
$\leq$ $\|X_{\lambda}^{x}-X_{\lambda}^{y}\|_{H}^{2}+\sup_{\lambda\leq\tau\leq\Lambda}\|X_{\tau}^{x}-X_{\tau}^{y}\|_{H}^{2}\int_{\lambda}^{\Lambda}C_{2}e^{\gamma(r-\Lambda)}\|X_{r}^{x}\|_{V}^{\alpha}\|X_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r.$
After taking the expected value, we get by the stochastic Gronwall’s Lemma [3, Lemma A.1], for $\kappa:=\frac{\varepsilon^{2}}{9\|F\|_{\operatorname{Lip}}^{2}}$, and for any $R>0$, by Proposition 5.3,
($\mathbb{P}\left(\left\{\|X_{t}^{x}-X_{t}^{y}\|^{2}_{H}\geq\frac{\varepsilon^{2}}{9\|F\|_{\operatorname{Lip}}^{2}}\right\}\right)$
$\leq$ ($\mathbb{P}\left(\left\{\sup_{0\leq r\leq t}\|X_{r}^{x}-X_{r}^{y}\|^{2}_{H}+(\delta_{2}-\gamma)\int_{0}^{t}\|X_{r}^{x}-X_{r}^{y}\|_{H}^{2}\,\mathrm{d}r\geq\kappa\right\}\right)$
$\leq$ $\frac{4}{\kappa}e^{4R}\|x-y\|_{H}^{2}+\mathbb{P}\left(\left\{\int_{0}^{t}C_{2}e^{\gamma(r-t)}\|X_{r}^{x}\|_{V}^{\alpha}\|X_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r\geq R\right\}\right)$
$\leq$ $\frac{4}{\kappa}e^{4R}\|x-y\|_{H}^{2}+\frac{1}{R}\mathbb{E}\left[\int_{0}^{t}C_{2}e^{\gamma(r-t)}\|X_{r}^{x}\|_{V}^{\alpha}\|X_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r\right]$
$\leq$ $\frac{36\|F\|^{2}_{\operatorname{Lip}}}{\varepsilon^{2}}e^{4R}\delta^{2}+\frac{2C_{2}}{\lambda_{0}\delta_{1}(\beta+2)R}\left(\|x\|_{H}^{\beta+2}+\frac{\mathfrak{c}}{\gamma}\right).$
First, choose $R>0$ such that the second term is less than $\frac{\varepsilon}{6\|F\|_{\infty}}$. Then, choose $\delta>0$ such that the first term is less than $\frac{\varepsilon}{6\|F\|_{\infty}}$. By (5.20), we get that,
$|P_{t}F(x)-P_{t}F(y)|<\frac{\varepsilon}{3}+2\|F\|_{\infty}\frac{2\varepsilon}{6\|F\|_{\infty}}=\varepsilon,$
whenever $\|x-y\|_{H}<\delta$. As $\delta>0$ can be chosen independent of $t\geq 0$, the claim is proved. ∎
Lemma 5.6.
Assume that Hypotheses (A)(B)(C)(D)(E), and (F) are valid. Then the semigroup $(P_{t})_{t\geq 0}$ associated to (2.1) is a stochastically continuous Markovian Feller semigroup.
Proof.
Let us prove the Feller property. Let $F\in C_{b}(H)$, $t\geq 0$. Hence, we get that,
$\|P_{t}F(\cdot)\|_{\infty}=\sup_{x\in H}|\mathbb{E}[F(X_{t}^{x})]|\leq\sup_{y\in H}\|F(y)\|_{H}<\infty.$
Let $x_{n},x\in H$, $n\in\mathbb{N}$ such that $\|x_{n}-x\|_{H}\to 0$ as $n\to\infty$. Let $t\geq 0$ and $F\in C_{b}(H)$. Let $F_{m}\in\operatorname{Lip}_{b}(H)$, $m\in\mathbb{N}$ with $\|F_{m}(\cdot)-F(\cdot)\|_{\infty}\to 0$ as $m\to\infty$. Then, for any $\varepsilon>0$, there exists $m\in\mathbb{N}$ such that
$\|F_{m}(\cdot)-F(\cdot)\|_{\infty}<\frac{\varepsilon}{3}.$
By the $e$-property proved in Proposition 5.5, for any $\delta>0$ there exists $n_{0}\in\mathbb{N}$ such that $\|x_{n}-x\|_{H}<\delta$ for all $n\geq n_{0}$, and thus
$|P_{t}F_{m}(x_{n})-P_{t}F_{m}(x)|<\frac{\varepsilon}{3}$
for $n\geq n_{0}(\delta)$ and all $t\geq 0$, whenever $\delta=\delta(\varepsilon,m)>0$ is small enough. As a consequence,
$|P_{t}F(x_{n})-P_{t}F(x)|$
$\leq$ $|P_{t}F(x_{n})-P_{t}F_{m}(x_{n})|+|P_{t}F_{m}(x_{n})-P_{t}F_{m}(x)|+|P_{t}F_{m}(x)-P_{t}F(x)|$
$=$ $|\mathbb{E}[F(X_{t}^{x_{n}})-F_{m}(X_{t}^{x_{n}})]|+|P_{t}F_{m}(x_{n})-P_{t}F_{m}(x)|+|\mathbb{E}[F_{m}(X_{t}^{x})-F(X_{t}^{x})]|$
$\leq$ $2\sup_{y\in H}\|F_{m}(y)-F(y)\|_{H}+\frac{\varepsilon}{3}=\varepsilon.$
Let us prove the stochastic continuity. Let $x\in H$. By [20, Proposition 2.1.1], it is sufficient to prove that
$|P_{t}F(x)-F(x)|\to 0$
as $t\searrow 0$ for any $F\in\operatorname{Lip}_{b}(H)$. Clearly,
$|P_{t}F(x)-F(x)|^{2}\leq\operatorname{Lip}(F)^{2}\mathbb{E}[\|X_{t}^{x}-x\|^{2}_{H}].$
By an adaptation of Corollary 5.4,
$\mathbb{E}\left[\|X_{t}^{x}-x\|^{\beta+2}_{H}\right]\leq\mathfrak{c}t\longrightarrow 0,$
as $t\searrow 0$.
The Markov property follows by the same arguments as in [56, Proposition 4.3.5], see also [34, Section 6.4]. ∎
5.3. Uniqueness of invariant measures
Our aim is to verify the conditions (3.9) and (3.10) of Theorem 3.10, compare with [34, 28].
For a bounded subset $J\subset H$, set $\|J\|_{H}:=\sup_{x\in J}\|x\|_{H}$. Set
$P_{t}(x,B):=P^{\ast}_{t}\delta_{x}(B),\quad t\geq 0,\quad x\in H,\quad B\in{\mathcal{B}}(H).$
Consider the measurable Lyapunov function
(5.21) $\Theta(x):=c^{\alpha}_{0}\delta_{1}\|x\|_{H}^{\alpha},$
which has bounded sublevel sets. Consider the locally monotone PDE
$\frac{\mathrm{d}}{\mathrm{d}t}u^{x}_{t}=A(u^{x}_{t}),\quad u^{x}_{0}=x\in H.$
The existence and uniqueness of solutions in $C([0,T];H)$ follows under Hypotheses (A)(B)(C)(D) by [56, Theorem 5.1.3].
Lemma 5.7.
Assume that Hypotheses (A)(B)(C)(D), and (F) hold true. For every $R>0$,
$\lim_{t\to\infty}\sup_{x\in B_{R}}\|u^{x}(t)\|_{H}=0,$
where $B_{R}:=\{y\in H:\Theta(y)\leq R\}$, where $\Theta$ is as in (5.21).
Proof.
Let $R>0$ and $x\in B_{R}$. By Hypothesis (B) and Fact 3.1 we have
$\frac{\mathrm{d}}{\mathrm{d}t}\|u_{t}^{x}\|_{H}^{2}=2\langle A(u^{x}_{t}),u^{x}_{t}\rangle\leq-\delta_{1}\|u_{t}^{x}\|^{\alpha}_{V}\leq-c^{\alpha}_{0}\delta_{1}\left(\|u_{t}^{x}\|_{H}^{2}\right)^{\alpha/2}.$
Note that $y(t):=\|u_{t}^{x}\|^{2}_{H}$ is a subsolution to the ordinary differential equation
$y^{\prime}=-c^{\alpha}_{0}\delta_{1}y^{\alpha/2},\quad\text{for a.e.}\;t\in[0,T],\quad y(0)=\|x\|^{2}_{H},$
with solution for $\alpha>2$,
$y(t)=\left(\|x\|^{2-\alpha}_{H}+\frac{\alpha-2}{2}c^{\alpha}_{0}\delta_{1}t\right)^{-2/(\alpha-2)},$
and solution for $\alpha=2$,
$y(t)=\|x\|_{H}^{2}\exp(-c_{0}^{2}\delta_{1}t).$
Hence, by a standard comparison principle, we have for $\alpha>2$ that
(5.22) $\begin{split}\|u_{t}^{x}\|_{H}&\leq\left(\|x\|^{2-\alpha}_{H}+\frac{\alpha-2}{2}c^{\alpha}_{0}\delta_{1}t\right)^{-1/(\alpha-2)}\\ &\leq\left(\left(\frac{R}{c_{0}^{\alpha}\delta_{1}}\right)^{(2-\alpha)/\alpha}+\frac{\alpha-2}{2}c^{\alpha}_{0}\delta_{1}t\right)^{-1/(\alpha-2)},\end{split}$
and for $\alpha=2$ that
$\|u_{t}^{x}\|_{H}\leq\left(\|x\|^{2}_{H}\exp(-c_{0}^{2}\delta_{1}t)\right)^{1/2}\leq\frac{\sqrt{R}}{c_{0}\sqrt{\delta_{1}}}\exp\left(-\frac{1}{2}c_{0}^{2}\delta_{1}t\right),$
which proves the claim. ∎
The next result shows uniform stochastic stability of the solutions to the SPDE and the deterministic PDE with positive probability.
Lemma 5.8.
Assume that Hypotheses (A)(B)(C)(D)(E)(F), and (G) hold true. For any $T>0$, for any $\varepsilon>0$, and $K\subset H$ bounded, such that we have
$\mathbb{P}\left(\|X_{T}^{x}-u_{T}^{x}\|_{H}^{2}<\varepsilon\right)>0,$
uniformly for $x\in K$.
Proof.
Let $T>0$, $\varepsilon>0$, $x\in K$. Note that $Y_{t}^{x}:=X_{t}^{x}-L_{t}$ is absolutely continuous and solves the non-autonomous random PDE
(5.23) $\frac{\mathrm{d}}{\mathrm{d}t}Y^{x}_{t}=A(Y^{x}_{t}+L_{t}),\quad Y_{0}=x,$
where $(L_{t})_{t\in[0,T]}$ is as in (3.4). By Lemma C.1, for any $\delta>0$, there exists an event $\Omega_{\delta}\in{\mathcal{F}}$ with positive probability such that
$\sup_{t\in[0,T]}\|L_{t}(\omega)\|_{V}<\frac{\delta}{2}$
for all $\omega\in\Omega_{\delta}$. Let us assume that $0<\delta<\sqrt{\varepsilon}$. Hence, for $\omega\in\Omega_{\delta}$, we have by Hypothesis (E) and the chain rule that
(5.24) $\begin{split}\|Y_{t}^{x}(\omega)\|_{H}^{2}=&\|x\|^{2}_{H}+2\int_{0}^{t}\langle A(Y_{s}^{x}(\omega)+L_{s}(\omega)),Y_{s}^{x}(\omega)\rangle\,\mathrm{d}s\\ =&\|x\|^{2}_{H}+2\int_{0}^{t}\langle A(X_{s}^{x}(\omega)),X_{s}^{x}(\omega)-L_{s}(\omega)\rangle\,\mathrm{d}s\\ \leq&\|x\|^{2}_{H}-\delta_{4}\int_{0}^{t}\|A(X_{s}^{x}(\omega))\|_{V^{\ast}}\,\mathrm{d}s-2\int_{0}^{t}\langle A(X_{s}^{x}(\omega)),L_{s}(\omega)\rangle\,\mathrm{d}s+(C_{4}\vee 0)t\\ \leq&\|x\|^{2}_{H}-\left(\delta_{4}-2\sup_{s\in[0,t]}\|L_{s}(\omega)\|_{V}\right)\int_{0}^{t}\|A(X^{x}_{s}(\omega))\|_{V^{\ast}}\,\mathrm{d}s+(C_{4}\vee 0)t\\ \leq&\|x\|_{H}^{2}-(\delta_{4}-\delta)\int_{0}^{t}\|A(X^{x}_{s}(\omega))\|_{V^{\ast}}\,\mathrm{d}s+(C_{4}\vee 0)t,\quad t\in[0,T].\end{split}$
We proceed similarly for $u^{x}_{t}$. We get that there exists a constant $\widetilde{C}(T)>0$, depending on $T$, $\delta_{4}$ and $C_{4}$, such that
(5.25) $\int_{0}^{T}\left(\|A(u_{s}^{x})\|_{V^{\ast}}+\|A(X^{x}_{s}(\omega))\|_{V^{\ast}}\right)\,\mathrm{d}s\leq\widetilde{C}(T)(1+\|x\|^{2}_{H}),$
for $0<\delta<\frac{\delta_{4}}{2}$. By the chain rule, Hypothesis (C),
$\|Y^{x}_{t}(\omega)-u_{t}^{x}\|^{2}_{H}=2\int_{0}^{t}\langle A(X^{x}_{s}(\omega))-A(u_{s}^{x}),X^{x}_{s}(\omega)-L_{s}(\omega)-u_{s}^{x}\rangle\,\mathrm{d}s$
and hence, for $\eta:=\frac{\delta}{2}\widetilde{C}(T)(1+\sup_{y\in K}\|y\|^{2}_{H})$,
$\sup_{s\leq t\leq T}\|Y^{x}_{t}(\omega)-u_{t}^{x}\|^{2}_{H}$ $\leq\|Y^{x}_{s}(\omega)-u_{s}^{x}\|^{2}_{H}$
$\quad+\sup_{s\leq t\leq T}\|Y^{x}_{t}(\omega)-u_{t}^{x}\|_{H}^{2}\int_{s}^{T}C_{2}\|u_{r}^{x}\|_{V}^{\alpha}\|u_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r$
$\quad+\frac{\delta}{2}\int_{s}^{T}\|A(X^{x}_{r}(\omega))-A(u^{x}_{r})\|_{V^{\ast}}\,\mathrm{d}r$
$\leq\left(\sup_{s\leq t\leq T}\|Y^{x}_{t}(\omega)-u_{t}^{x}\|_{H}^{2}+\eta\right)\int_{s}^{T}C_{2}\|u_{r}^{x}\|_{V}^{\alpha}\|u_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r+\eta.$
Let $\mathbb{E}_{\delta}:=\mathbb{E}_{\mathbb{P}(\cdot\,|\,\Omega_{\delta})}$ be the expected value with respect to the conditional probability $\mathbb{P}(\cdot\,|\,\Omega_{\delta})$. Let $0\leq\lambda\leq\Lambda\leq T$ be stopping times. We get that
$\mathbb{E}_{\delta}\left[\sup_{\lambda\leq\tau\leq\Lambda}\|Y^{x}_{\tau}-u_{\tau}^{x}\|^{2}_{H}\right]$
$\leq$ $\left(\mathbb{E}_{\delta}\left[\|Y^{x}_{\lambda}-u_{\lambda}^{x}\|^{2}_{H}\right]+\eta\right)+\mathbb{E}_{\delta}\left[\left(\sup_{\lambda\leq\tau\leq\Lambda}\|Y^{x}_{\tau}-u_{\tau}^{x}\|_{H}^{2}+\eta\right)\int_{\lambda}^{\Lambda}C_{2}\|u_{r}^{x}\|_{V}^{\alpha}\|u_{r}^{x}\|_{H}^{\beta}\,\mathrm{d}r\right].$
Note that by the chain rule and Hypothesis (B) for $0\leq s\leq t\leq T$,
$\|u_{t}^{x}\|^{\beta+2}_{H}+\delta_{1}\frac{\beta+2}{2}\int_{s}^{t}\|u_{r}^{x}\|^{\alpha}_{V}\|u_{r}^{x}\|^{\beta}_{H}\,\mathrm{d}r\leq\|u_{s}^{x}\|^{\beta+2}_{H}.$
By the stochastic Gronwall Lemma [3, Lemma A.1],
($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}\geq\frac{\varepsilon}{4}\right\}\,\Bigg{|}\,\Omega_{\delta}\right)$
$\leq$ $\frac{16}{\varepsilon}e^{4R}\eta+\mathbb{P}\left(\left\{C_{2}\int_{0}^{T}\|u_{t}^{x}\|_{V}^{\alpha}\|u_{t}^{x}\|_{H}^{\beta}\,\mathrm{d}t\geq R\right\}\,\bigg{|}\,\Omega_{\delta}\right)$
$\leq$ $\frac{16}{\varepsilon}e^{4R}\eta+\frac{1}{R}\mathbb{E}_{\delta}\left[1_{\Omega_{\delta}}C_{2}\int_{0}^{T}\|u_{t}^{x}\|_{V}^{\alpha}\|u_{t}^{x}\|_{H}^{\beta}\,\mathrm{d}t\right]$
$\leq$ $\frac{8}{\varepsilon}e^{4R}\delta\widetilde{C}(T)\left(1+\sup_{y\in K}\|y\|^{2}_{H}\right)+\frac{1}{R}\frac{2C_{2}}{\delta_{1}(\beta+2)}\sup_{y\in K}\|y\|^{\beta+2}_{H}.$
First, choose $R>0$ such second term is smaller than $\frac{1}{4}$. Then choose $0<\delta<\sqrt{\varepsilon}\wedge\frac{\delta_{4}}{2}$ such that the first term is smaller than $\frac{1}{4}$. We get that
($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}\geq\frac{\varepsilon}{4}\right\}\,\Bigg{|}\,\Omega_{\delta}\right)<\frac{1}{2}.$
Hence
($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\,\Bigg{|}\,\Omega_{\delta}\right)\geq\frac{1}{2},$
uniformly for $x\in K$. And thus, by $\delta<\sqrt{\varepsilon}$,
($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|X^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\varepsilon\right\}\right)$
$\geq$ ($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\cap\left\{\sup_{0\leq t\leq T}\|L_{t}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\right)$
$=$ ($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\cap\Omega_{\sqrt{\varepsilon}}\right)$
$\geq$ ($\mathbb{P}\left(\left\{\sup_{0\leq t\leq T}\|Y^{x}_{t}-u_{t}^{x}\|^{2}_{H}<\frac{\varepsilon}{4}\right\}\cap\Omega_{\delta}\right)\geq\frac{\mathbb{P}(\Omega_{\delta})}{2}>0,$
uniformly for $x\in K$. The claim is proved. ∎
The next lemma verifies the second condition (3.10) in the lower bound technique in Theorem 3.10. Note that we do not need to verify tightness, as the method proves a powerful alternative the the Krylov-Bogoliubov method [20].
Lemma 5.9.
Assume that Hypotheses (A)(B)(C)(D), and (F) are valid. Then for each $\varepsilon>0$ and each bounded set $J\subset H$ there exists a constant $R(\varepsilon,\|J\|_{H})>0$ such that the sublevel set
$K:=\{\Theta(\cdot)\leq R(\varepsilon,\|J\|_{H})\}\subset H$
satisfies
$\inf_{x\in J}\liminf_{T\to\infty}Q_{T}(x,K)>1-\varepsilon.$
Here, $\Theta$ is as in (5.21).
Proof.
Let $\varepsilon>0$, $J\subset H$ be bounded and $x\in J$. For $R>0$, $K_{R}:=\{\Theta\leq R\}$ is a measurable set. By
$\langle A(x),x\rangle\geq-\Theta(x),$
and by Itô’s formula, assuming Hypotheses (B) and (F), there exists a constant $C>0$ such that
$\frac{1}{T}\mathbb{E}\int_{0}^{T}\Theta(X_{s}^{x})\,\mathrm{d}s\leq C\left(\|x\|_{H}^{2}+1\right),\quad\text{for}\;\;T\geq 1,$
compare also with (5.18). As a consequence,
$Q_{T}(x,K_{R})=$ $\int_{0}^{T}P_{s}(x,K_{R})\,\mathrm{d}s\geq\frac{1}{T}\int_{0}^{T}\left(1-\frac{\mathbb{E}[\Theta(X_{s}^{x})]}{R}\right)\,\mathrm{d}s$
$\geq$ $1-\frac{C}{R}\left(\|J\|_{H}^{2}+1\right).$
Choosing $R(\varepsilon,\|J\|_{H})>\varepsilon^{-1}C(\|J\|^{2}_{H}+1)$ yields the claim with
$K:=K_{R(\varepsilon,\|J\|_{H})}=\{x\in H\,\colon\,\Theta(x)\leq R(\varepsilon,\|J\|_{H})\}.$
The next lemma verifies the first condition (3.9) in the lower bound technique in Theorem 3.10.
Lemma 5.10.
Assume Hypotheses (A)(B)(C)(D)(E)(F), and (G). Then for each $\delta>0$ and each bounded set $J\subset H$,
$\inf_{x\in J}\liminf_{T\to\infty}Q_{T}(x,B_{\delta}(0))>0.$
Proof.
Let $\delta>0$, $J\subset H$ be bounded, $x\in J$ and $K=K_{R(\frac{1}{2},\|J\|_{H})}$ be as in Lemma 5.9. By Lemma 5.7, there exists $T_{0}>0$ corresponding to $K$ such that
$\|u_{T_{0}}^{z}\|_{H}\leq\frac{\delta}{2}$
for all $z\in K$. Using Lemma 5.8, yields
$P_{T_{0}}(z,B_{\delta}(0))=\mathbb{P}(\|X^{z}_{T_{0}}\|_{H}\leq\delta)\geq\mathbb{P}\left(\|X^{z}_{T_{0}}-u^{z}_{T_{0}}\|_{H}\leq\frac{\delta}{2}\right)\geq\gamma_{1}>0,$
where $\gamma_{1}=\gamma_{1}(T_{0},\delta)$ is independent of $z\in K$. Thus by Lemma 5.9 with $\varepsilon=\frac{1}{2}$,
$\liminf_{T\to\infty}Q_{T}(x,B_{\delta}(0))=\liminf_{T\to\infty}\frac{1}{T}\int_{0}^{T}P_{s}(x,B_{\delta}(0))\,\mathrm{d}s$
$=$ $\liminf_{T\to\infty}\frac{1}{T}\int_{0}^{T}P_{s+T_{0}}(x,B_{\delta}(0))\,\mathrm{d}s=\liminf_{T\to\infty}\frac{1}{T}\int_{0}^{T}\int_{H}P_{s}(x,\mathrm{d}z)P_{T_{0}}(z,B_{\delta}(0))\,\mathrm{d}s$
$\geq$ $\liminf_{T\to\infty}\frac{1}{T}\int_{0}^{T}\int_{K}P_{s}(x,\mathrm{d}z)P_{T_{0}}(z,B_{\delta}(0))\,\mathrm{d}s\geq\liminf_{T\to\infty}\gamma_{1}\frac{1}{T}\int_{0}^{T}\int_{K}P_{s}(x,\mathrm{d}z)\,\mathrm{d}s$
$\geq$ $\gamma_{1}\liminf_{T\to\infty}Q_{T}(x,K)\geq\frac{\gamma_{1}}{2}>0,$
where $\gamma_{1}=\gamma_{1}(T_{0},\delta)=\gamma_{1}(\|J\|_{H},\delta)$. ∎
Note that for proving the previous lemma, the estimate (5.2) would not be sufficient, as we cannot guarantee that small sublevels have lower bounded transition probabilities uniformly on bounded sets for large times. Therefore, we need to use the coupling with the deterministic flow, which has a good decay behavior by Lemma 5.7.
Thus we have proved the existence of a unique invariant measure $\mu_{\ast}$ for $(P_{t})_{t\geq 0}$ and thus the first part of Theorem 2.1 by an application of Theorem 3.10. The concentration property can be seen as follows:
Lemma 5.11 (Concentration).
Assume that Hypotheses (A)(B)(C)(D) and (F) are valid. Then any invariant probability measure $\mu_{\ast}$ has finite $H$-moments of order $\alpha+\beta$. In addition,
(5.26) $\int_{H}\|z\|^{\alpha+\beta}_{H}\,\mu_{\ast}(\mathrm{d}z)\leq\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}.$
In particular,
(5.27) $\int_{H}\|z\|_{H}\,\mu_{\ast}(\mathrm{d}z)\leq\left(\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\right)^{1/(\alpha+\beta)}.$
If $\beta=0$, $\mu_{\ast}$ has finite $V$-moments of order $\alpha$, and, in particular,
(5.28) $\int_{H}\|z\|^{\alpha}_{V}\,\mu_{\ast}(\mathrm{d}z)\leq\frac{C_{2}}{\lambda_{0}\delta_{1}\gamma}\mathfrak{c},$
and
(5.29) $\int_{H}\|z\|_{V}\,\mu_{\ast}(\mathrm{d}z)\leq\left(\frac{C_{2}}{\lambda_{0}\delta_{1}\gamma}\mathfrak{c}\right)^{1/\alpha}.$
Here, $\gamma\in(0,\delta_{2})$ and $\mathfrak{c}=\sum_{i=1}^{5}c_{i}\geq 0$ are as in Definition 5.1.
Proof.
By (5.18) and Fact 3.1, it follows for every $z\in H$, for every $t\geq 0$, and every $\ell\in\mathbb{N}$ that
(5.30) $\begin{split}\left(\mathbb{E}\left[\int_{0}^{t}\|X_{s}^{z}\|_{H}^{\alpha+\beta}\,\mathrm{d}s\right]\right)\wedge\ell&\leq\left(\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}t\right)\wedge\ell+\left(\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\|z\|_{H}^{\beta+2}\right)\wedge\ell.\end{split}$
Since $\mu_{\ast}$ is invariant, we have
(5.31) $\begin{split}\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)=\int_{H}\mathbb{E}\left[\|X^{z}_{t}\|^{\alpha+\beta}_{H}\wedge\ell\right]\,\mu_{\ast}(\mathrm{d}z).\end{split}$
With the help of Fubini’s Theorem, the preceding equality yields for $t>0$,
(5.32) $\begin{split}\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\mu_{\ast}(\mathrm{d}z)&=\frac{1}{t}\int_{0}^{t}\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)\,\mathrm{d}s\\ &=\frac{1}{t}\int_{0}^{t}\int_{H}\mathbb{E}\left[\|X^{z}_{s}\|^{\alpha+\beta}_{H}\wedge\ell\right]\,\mu_{\ast}(\mathrm{d}z)\,\mathrm{d}s\\ &=\int_{H}\mathbb{E}\left[\frac{1}{t}\int_{0}^{t}\,\left(\|X^{z}_{s}\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mathrm{d}s\right]\,\mu_{\ast}(\mathrm{d}z)\\ &\leq\int_{H}\left(\left(\frac{1}{t}\int_{0}^{t}\mathbb{E}\left[\|X^{z}_{s}\|^{\alpha+\beta}_{H}\right]\,\mathrm{d}s\right)\wedge\ell\right)\mu_{\ast}(\mathrm{d}z).\end{split}$
The preceding inequality, together with (5.30), implies for $\alpha>2$, $t>0$,
$\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\int_{H}\left(\|z\|^{\beta+2}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\int_{H}\left(\left(\|z\|^{\alpha+\beta}_{H}+1\right)\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z),$
and thus for $t\geq\frac{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}{2C_{2}}$,
$\left(1-\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\right)\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}(\beta+2)}\wedge\ell.$
For $\alpha=2$, we get for $t>0$ that
$\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\delta_{1}(\beta+2)}\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z),$
and thus $t\geq\frac{c_{0}^{\alpha}\delta_{1}(\beta+2)}{2C_{2}}$,
$\left(1-\frac{1}{t}\frac{2C_{2}}{c_{0}^{\alpha}\delta_{1}(\beta+2)}\right)\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\,\mu_{\ast}(\mathrm{d}z)$
$\leq$ $\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell.$
For $\alpha\geq 2$, and for any $\varepsilon>0$, we can find a large enough $t>0$, such that
(5.33) $\int_{H}\left(\|z\|^{\alpha+\beta}_{H}\wedge\ell\right)\mu_{\ast}(\mathrm{d}z)\leq\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}\wedge\ell+\varepsilon.$
After sending $\ell\to\infty$, we obtain with the help of Fatou’s lemma that
(5.34) $\int_{H}\|z\|^{\alpha+\beta}_{H}\mu_{\ast}(\mathrm{d}z)\leq\frac{2C_{2}}{c_{0}^{\alpha}\lambda_{0}\delta_{1}\gamma(\beta+2)}\mathfrak{c}+\varepsilon.$
Inequality (5.26) follows from the fact that $\varepsilon>0$ was arbitrary. Inequality (5.27) follows directly by Jensen’s inequality.
For $\beta=0$, (5.18) implies for $\ell\in\mathbb{N}$, $z\in H$, $t\geq 0$,
(5.35) $\begin{split}\left(\mathbb{E}\left[\int_{0}^{t}\|X_{s}^{z}\|_{V}^{\alpha}\,\mathrm{d}s\right]\right)\wedge\ell&\leq\left(\frac{C_{2}}{\lambda_{0}\delta_{1}\gamma}\mathfrak{c}t\right)\wedge\ell+\left(\frac{C_{2}}{\lambda_{0}\delta_{1}}\|z\|_{H}^{\beta+2}\right)\wedge\ell.\end{split}$
and hence the proof can be completed by repeating the argument above. ∎
Appendix A An estimate in Hilbert spaces
The following inequality was noted in [14, Equation (4.9)] without the explicit constant for $\beta>0$.
Lemma A.1.
In any Hilbert space $H$, we have any $\beta\geq 0$ and for
$C_{\beta}:=((2^{\beta-3}\vee 1)\beta(\beta+2))\vee((2^{\beta-1}\vee 1)(\beta+2)),\quad\beta>0,$
and $C_{0}:=\frac{1}{2}$ that
$0\leq\|x+h\|_{H}^{\beta+2}-\|x\|^{\beta+2}_{H}-(\beta+2)\|x\|^{\beta}_{H}\left\langle x,h\right\rangle\leq C_{\beta}\left(\|x\|_{H}^{\beta}\|h\|_{H}^{2}+\|h\|^{\beta+2}_{H}\right),$
for any $x,h\in H$, where the second inequality is an equality for $\beta=0$ and $C_{0}=\frac{1}{2}$.
Proof.
The first inequality follows from the convex subdifferential inequality, see e.g. [24], noting that the first variation of the functional $F:x\mapsto\|x\|_{H}^{\beta+2}$ is
$\partial F(x)(h)=(\beta+2)\|x\|^{\beta}_{H}\left\langle x,h\right\rangle.$
The second variation of $F$ is given by
$\partial^{2}F(x)(h,h)=\beta(\beta+2)\|x\|^{\beta-2}|\langle x,h\rangle|^{2}+(\beta+2)\|x\|^{\beta}\|h\|^{2}\quad\text{if }x\not=0,$
and $\partial^{2}F(x)(h,h)=0$ if $x=0$.
By Taylor’s formula for Fréchet differentials, see e.g. [79],
$|\|x+h\|_{H}^{\beta+2}-\|x\|^{\beta+2}_{H}-(\beta+2)\|x\|^{\beta}_{H}\left\langle x,h\right\rangle|\leq\frac{1}{2}\partial^{2}F(\xi_{\lambda})(h,h)$
for some $\xi_{\lambda}=\lambda x+(1-\lambda)h$, $\lambda\in(0,1)$. Noting that $\|\xi_{\lambda}\|_{H}\leq\|x\|_{H}+\|h\|_{H}$, we get the claim. The equality for $\beta=0$ follows follows from the polarization identity in Hilbert spaces. ∎
Appendix B A convexity result
For $\alpha\geq 2$ and $\beta\geq 0$, define the function $g:\mathbb{R}^{2}\to\mathbb{R}$ by
(B.1) $g:(x,y)\mapsto|x|^{\alpha}|y|^{\beta}.$
Lemma B.1.
For $\alpha\geq 2$ and $\beta\geq 0$, $g$ as defined in (B.1) is convex.
Proof.
The gradient of $g$ is given for $y\not=0$ by
$\nabla g(x,y)=\begin{pmatrix}\alpha|x|^{\alpha-2}x|y|^{\beta}\\ \beta|x|^{\alpha}|y|^{\beta-2}y\end{pmatrix},$
and
$\nabla g(x,0)=(0,0)^{\textup{t}}.$
The Hessian matrix is given for $y\not=0$ by
$\operatorname{Hess}g(x,y)=\begin{pmatrix}(\alpha-2)\alpha|x|^{\alpha-2}|y|^{\beta}+\alpha|x|^{\alpha-2}&\alpha\beta xy|x|^{\alpha-2}|y|^{\beta-2}\\ (\beta-2)\beta|x|^{\alpha}|y|^{\beta-2}+\beta|x|^{\alpha}|y|^{\beta-2}&(\beta-2)\beta|x|^{\alpha}|y|^{\beta-2}\end{pmatrix},$
and
$\operatorname{Hess}g(x,0)=\begin{pmatrix}\alpha|x|^{\alpha-2}&0\\ 0&0\end{pmatrix}.$
We get that the determinant is non-negative,
$\det\operatorname{Hess}g(x,y)=|x|^{\alpha}|y|^{\beta}\geq 0,$
as well as the first entry is non-negative,
$(\alpha-2)\alpha|x|^{\alpha-2}|y|^{\beta}+\alpha|x|^{\alpha-2}\geq 0,$
for $\alpha\geq 2$. Hence, by Sylvester’s criterion, $\operatorname{Hess}g(x,y)$ is non-negative definite for every $x,y\in\mathbb{R}$, which implies that $g$ is convex. ∎
Lemma B.2.
For $x,y\in V$, $\alpha\geq 2$, $\beta\geq 0$, the map $\Phi:V\times V\to\mathbb{R}$ defined by $\Phi:(x,y)\mapsto\|x\|^{\alpha}_{V}\|y\|^{\beta}_{H}$ is convex and lower semi-continuous with respect to weak convergence in $V\times V$.
Proof.
Let $\lambda\in[0,1]$, $x_{1},x_{2},y_{1},y_{2}\in V$. As norms are convex, by Lemma B.1, for $g$ as in (B.1), noting that both components of $g$ are increasing,
$\Phi(\lambda x_{1}+(1-\lambda)x_{2},\lambda y_{1}+(1-\lambda)y_{2})$
$=$ $g(\|\lambda x_{1}+(1-\lambda)x_{2}\|_{V},\|\lambda y_{1}+(1-\lambda)y_{2}\|_{H})$
$\leq$ $g(\lambda\|x_{1}\|_{V}+(1-\lambda)\|x_{2}\|_{V},\lambda\|y_{1}\|_{H}+(1-\lambda)\|y_{2}\|_{H})$
$\leq$ $\lambda g(\|x_{1}\|_{V},\|y_{1}\|_{H})+(1-\lambda)g(\|x_{2}\|_{V},\|y_{2}\|_{H})$
$=$ $\lambda\Phi(x_{1},y_{1})+(1-\lambda)\Phi(x_{2},y_{2}).$
As $\Phi$ is obviously strongly lower semi-continuous in $V\times V$ by the continuous embedding $V\subset H$, we get by Mazur’s lemma, see e.g. [24], that $\Phi$ is weakly lower semi-continuous in $V\times V$. ∎
Appendix C Small ball probability
We shall need a result on the so-called small ball problem for Lévy processes. As in Hypothesis (G), set
(C.1) $L_{t}:=B\,W_{t}+\int_{D^{c}}G(w)\,\widetilde{N}(t,\mathrm{d}w),\quad t\in[0,T],$
where $(W_{t})_{t\in[0,T]}$ and $\left(\int_{D^{c}}w\,\widetilde{N}(t,\mathrm{d}w)\right)_{t\in[0,T]}$ are independent.
Lemma C.1.
Under Hypotheses (F) and (G), for any $\delta>0$,
(C.2) $\mathbb{P}\left(\sup_{t\in[0,T]}\left\|L_{t}\right\|_{V}<\delta\right)>0,$
where $(L_{t})_{t\in[0,T]}$ is as in (3.4).
Proof.
Fix $T>0$ and $\delta>0$. Let $(L_{t})_{t\in[0,T]}$ have the representation as in Hypothesis (G). For $N\in\mathbb{N}$ denote
$L_{t}^{N}:=\sum_{k=1}^{N}\sqrt{\lambda_{k}}e_{k}l_{t}^{k},$
and let $\lceil T\rceil$ be the smallest integer greater or equal to $T$. By infinite divisibility of $(l^{k})_{s\in[0,T]}$, there exist $\lceil T\rceil$ i.i.d. copies of symmetric Lévy processes $(l^{k,i}_{s})_{s\in\left[0,\frac{T}{\lceil T\rceil}\right]}$, $i=1,\ldots,\lceil T\rceil$ such that
$l_{s}^{k}=\sum_{i=1}^{\lceil T\rceil}l_{\frac{s}{\lceil T\rceil}}^{k,i},\quad s\in[0,T].$
By independence,
$\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}^{N}\|_{V}<\frac{\delta}{2}\right)=\mathbb{P}\left(\sup_{t\in[0,T]}\left\|\sum_{k=1}^{N}\sqrt{\lambda_{k}}e_{k}l_{t}^{k}\right\|_{V}<\frac{\delta}{2}\right)$
$\geq$ $\mathbb{P}\left(\sup_{t\in[0,T]}\sum_{k=1}^{N}\sqrt{\lambda_{k}}|l_{t}^{k}|\|e_{k}\|_{V}<\frac{\delta}{2}\right)\geq\mathbb{P}\left(\sup_{t\in[0,T]}\sum_{k=1}^{N}\sqrt{\lambda_{k}\sigma_{k}}|l_{t}^{k}|<\frac{\delta}{2}\right)$
$\geq$ $\mathbb{P}\left(\sup_{1\leq k\leq N}\sup_{t\in[0,T]}\left(\sqrt{\lambda_{k}\sigma_{k}}|l_{t}^{k}|\right)<\frac{\delta}{2N}\right)\geq\mathbb{P}\left(\bigcap_{k=1}^{N}\left\{\sup_{t\in[0,T]}\left(\sqrt{\lambda_{k}\sigma_{k}}|l_{t}^{k}|\right)<\frac{\delta}{2N}\right\}\right)$
$=$ $\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,T]}\left(\sqrt{\lambda_{k}\sigma_{k}}|l_{t}^{k}|\right)<\frac{\delta}{2N}\right).$
By [7, Corollary 2.1], there exists $K>0$ with
$\begin{split}&-\log\mathbb{P}\left(\sup_{t\in[0,1]}|l_{t}^{k,1}|<\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}}\right)\\ =&K\left[\tilde{\pi}(I^{c}_{k,N,\delta,T})+\frac{4N^{2}\lceil T\rceil^{2}\lambda_{k}\sigma_{k}}{\delta^{2}}\left(\tilde{\sigma}^{2}+\int_{I_{k,N,\delta,T}}u^{2}\,\tilde{\pi}(\mathrm{d}u)\right)\right]=:Ka_{k,N,\delta,T},\end{split}$
where $I_{k,N,\delta,T}:=\left[-\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}},\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}}\right]$, and $(\tilde{\sigma}^{2},\tilde{\pi},0)$ is the characteristic triple of the symmetric Lévy process $(l_{t}^{k,1})_{t\in[0,1]}$, which does not depend on $k$. By Hypothesis (F) we have
$\tilde{\pi}(I_{k,N,\delta,T}^{c})<\infty,\quad\int_{I_{k,N,\delta,T}}u^{2}\,\tilde{\pi}(\mathrm{d}u)<\infty,$
and then,
$\begin{split}&\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,T]}|l_{t}^{k}|<\frac{\delta}{2N\sqrt{\lambda_{k}\sigma_{k}}}\right)=\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,T]}\left|\sum_{i=1}^{\lceil T\rceil}l_{t/\lceil T\rceil}^{k,i}\right|<\frac{\delta}{2N\sqrt{\lambda_{k}\sigma_{k}}}\right)\\ \geq&\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,T/\lceil T\rceil]}|l_{t}^{k,1}|<\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}}\right)\geq\prod_{k=1}^{N}\mathbb{P}\left(\sup_{t\in[0,1]}|l_{t}^{k,1}|<\frac{\delta}{2N\lceil T\rceil\sqrt{\lambda_{k}\sigma_{k}}}\right)\\ =&\exp\left(-K\sum_{k=1}^{N}a_{k,N,\delta,T}\right)\in(0,1),\end{split}$
compare also [43, Lemma 5.6].
By Doob’s submartingale inequality and Hypothesis (G),
$\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}-L_{t}^{N}\|_{V}<\frac{\delta}{2}\right)=\mathbb{P}\left(\sup_{t\in[0,T]}\left\|\sum_{k=N+1}^{\infty}\sqrt{\lambda_{k}}e_{k}l_{t}^{k}\right\|_{V}^{2}<\frac{\delta^{2}}{4}\right)$
$\geq$ $\mathbb{P}\left(\sup_{t\in[0,T]}\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}|l_{t}^{k}|^{2}<\frac{\delta^{2}}{4}\right)=1-\mathbb{P}\left(\left\{\exists t_{0}\in[0,T]\;\colon\;\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}|l_{t_{0}}^{k}|^{2}\geq\frac{\delta^{2}}{4}\right\}\right)$
$\geq$ $1-\mathbb{P}\left(\sup_{t\in[0,T]}\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}|l_{t}^{k}|^{2}\geq\frac{\delta^{2}}{4}\right)\geq 1-\frac{4}{\delta^{2}}\mathbb{E}[|l^{1}_{T}|^{2}]\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}.$
Observe that
$01-\frac{4}{\delta^{2}}\mathbb{E}[|l^{1}_{T}|^{2}]\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}>$ 0
if and only if
$\frac{4}{\delta^{2}}\mathbb{E}[|l^{1}_{T}|^{2}]\sum_{k=N+1}^{\infty}\lambda_{k}\sigma_{k}<1.$
Recalling that $\{\lambda_{k}\sigma_{k}\}_{k\in\mathbb{N}}\in\ell^{1}$ by Hypothesis (G), for large enough $N\in\mathbb{N}$,
$\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}(\omega)-L_{t}^{N}(\omega)\|_{V}<\frac{\delta}{2}\right)>0.$
Hence, by independence,
$\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}\|_{V}<\delta\right)\geq\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}-L_{t}^{N}\|_{V}<\frac{\delta}{2}\right)\cdot\mathbb{P}\left(\sup_{t\in[0,T]}\|L_{t}^{N}\|_{V}<\frac{\delta}{2}\right)>0.$
Declarations
Acknowledgments. GB would like to express his gratitude to University of Helsinki, Aalto University School of Science, and Instituto Superior Técnico for all the facilities used along the realization of this work. JMT would like to thank the Department of Mathematics and Systems Analysis, Aalto University for providing its facilities and free coffee for this research. JMT would like to thank Erika Hausenblas (Department of Mathematics, Montanuniversität Leoben, Austria) for providing the important references [7, 43]. JMT would like to thank Mark Veraar (Delft Institute of Applied Mathematics, Delft University of Technology, Netherlands) for an inspiring discussion and for suggesting to apply the stochastic Gronwall inequality from [32]. JMT would like to thank Dirk Blömker (Institut für Mathematik, Universität Augsburg, Germany) for fruitful discussions on stability of the 2D stochastic Navier-Stokes equations.
Funding. The research of GB has been supported by the Finnish Centre of Excellence in Randomness and Structures (decision numbers 346306 and 346308). The research of both authors was partially supported by Horizon Europe Marie Skłodowska-Curie Actions Staff Exchanges (project no. 101183168 – LiBERA).
Ethical approval. Not applicable.
Competing interests. The authors declare that they have no conflict of interest.
Authors’ contributions. All authors have contributed equally to the paper.
Availability of data and materials. Data sharing not applicable to this article as no data-sets were generated or analyzed during the current study.